A Complete Bibliography of The American Statistician: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

07 October 2024 Version 1.04

Title word cross-reference

2 × 2 [Gri67]. b_{0} [SG69]. b_{1} [SG69]. R̄^{2} [Edw69]. β [Cur64]. χ^{2} [Gri67]. F [AR67, Bai69, Cur64, Duc69, Edw69, Lau65]. k [JH67, Mey69]. n [Mey69]. n = 3 [JR68]. r [JH67]. R/X [Ano68l, Kir68]. R^{2} [Hai69]. σ^{p} [Stu69]. t [And61, Bai69, Cur64, Duc69]. X [SG69]. Y [SG69]. y = ax [Ker66]. Z [Cur64].

-Distribution [Bai69]. -Tables [Lau65]. -Test [Gri67, Edw69]. -th [Mey69]. -Tuple [Mey69].

11 [Wer66]. 120th [Ano60-30]. 121st [Ano61-36]. 122nd [Ano62-42]. 123rd [Ano63-37]. 124th [Ano64a]. 125th [Ano65a, Ano65b]. 126th [Ano66a]. 127th [Ano67a]. 128th [Ano68a]. 129th [Ano69a]. 1960 [Ano60a, Hau60, Mar60, Pab60, Ste60b]. 1961 [Ano61k]. 1963 [Ano64u, Tae63]. 1964 [Ric64, Ros64]. 1965 [Ano65n]. 1966 [Ano66n, Mor66]. 1967 [Ano67b, Ano67s, Ano67z, Dem67]. 1968 [Ano68m, Cla68, SK68, Wor68]. 1969 [Eck69, GT69, Lew69a, Lew69b, Sib69].

26 [Ano60a].

A.S.A. [Van68]. Absolute [Beh67, Sem64, Sha69]. Abstract [Col65]. Access [KHH^{+}69]. Accurate [Shi60]. Acknowledgement [Mar68b]. Act [Bur62]. Actions [Mos67]. Activities [Ano60o, Ano60p, Ano60q, Ano60r, Ano60s, Ano61l, Ano61m, Ano61s, Ano61t, Ano61u, Ano61v, Ano61w, Ano61x, Ano62k, Ano62l, Ano62m, Ano62n, Ano62o, Ano62u, Ano62v, Ano62w, Ano62x, Ano62y, Ano62z, Ano62-27, Ano62-28, Ano63l, Ano63m, Ano63n, Ano63o, Ano63u, Ano63v, Ano63w, Ano63x, Ano63y, Ano63z, Ano64k, Ano64l, Ano64m, Ano64n, Ano64t, Ano64v, Ano64w, Ano64x, Ano64y, Ano65p, Ano65q, Ano65r, Ano65s, Ano65t, Ano65z, Ano65-28, Ano65-29, Ano65-30, Ano65-31, Ano65-32, Ano66p, Ano66q, Ano66r, Ano66s, Ano66y, Ano66z, Ano66-27, Ano66-28, Ano66-29, Ano66-30, Ano66-44, Ano67u, Ano67v, Ano67w, Ano67x, Ano67y, Ano68t, Ano68u, Ano68v, Ano68w, Ano69v, BTGK60, BKH61, KCSP61, LCD60, MWR60, RK60, Smi63, SBFC60, WGW^{+}61]. Activities [WDC^{+}60]. Adding [Hor60]. Additional [Li64]. Adjustment [Sie68]. Adventures [Mei60]. Advisory [Ano69o, SBB^{+}65]. Age [Ano66-41, Kar64]. Agencies [Ano66o, Ano69n, DS64]. Agricultural [Bea67a, HB67]. Aid [Nat60, War69]. Air [Har62]. Ajgaonkar [Han68]. Algebraic [SG69]. Alternate [Jae66]. Alternative [Mit69]. Alternatives [Che68]. American [Ano60-30, Ano61-36, Ano65-45, Ano66-45, Ano69-34, Jaf60, Ano62-40, Ano63k, Ano64u, BH68, RBB65, SBB^{+}65, Wal65a]. Analogue [Els69]. Analysis [And61, Arm67, Bro63, Cor66, Edg63, Hah68, Men68, Mit69, Pri67, SS66, Smi64a, War69, Ano65-44, Ano68-34]. Analytical [Har65]. Andrew [Hau60]. Anniversaries [Eis65, Eis67, EB67]. Anniversary [Ano65a, Ano69o]. Annual [Ano60a, Ano60-30, Ano61-36, Ano62-42, Ano63-37, Ano64a, Ano65b, Ano66a, Ano67b, Ano67a, Ano68a, Ano69a]. Answers [Ano66m, Ano67m, Rub60b, Rub60c, RS60, Rub60d, Rub61, RE61, RR61, RH61, RS61, Rub62a, Rub62b, Rub62c, Rub62d, Rub62e, Rub63a, Rub63b, RPG^{+}63, RCF63, Rub63c, Rub64a, Rub64b, Rub64c, Rub64d, Rub65a, Rub65b, RR65, Rub65c, Rub65d, RH66, RA66, RNWR66, Rub66a, Rub66b, RNCE67, Rub67a, RR67, Rub67b, RM67, Rub68a, Rub68b, Rub68c, Rub68d, Rub68e, Rub69a, Rub69b, Rub69c, RG69, Rub69d]. Ante [BM64]. Anticipations [Fer66, Fer67, Kat67]. Appeal [oH69]. Application [Ano66-44, Ano68l, Cox68, Kir68, Kul68]. Applications [Bak66, Bak67, LS61]. Applied [GS65, Lun68]. Approach [Jae66, Mit69, Sie68]. Approximate [Smi69]. Approximately [Sea67]. Approximation [Hoy68]. Approximations [BHW68]. Archival [oH69]. Area [BGB68, Bla68]. Army [Mal62]. Arthur [Sib69]. ASA [Ano62-29, Ano62-41, Ano64-34, Ano65c, Ano66b, Ano66-41, Ano66-42, Ano66-43, Ano66-44, GMN66, Mos67, Zar69]. Association [Ano65-45, Ano66-45, Ano69-34, Jaf60, Ano60-30, Ano61-36, Ano62-40, Ano63k, RBB65, SBB^{+}65, Wal65a]. ASTI [Lun68]. ASTM [Wer66]. Attainments [Ano66-42]. August [Ano60a, Ano64u]. Automation [LS61]. Averages [Man69a].

B [Ano65-27]. Back [Ano60b, Ano60c, Ano60d, Ano61a, Ano61b, Ano61c, Ano61d, Ano61e, Ano62a, Ano62b, Ano62c, Ano62d, Ano62e, Ano63a, Ano63b, Ano63c, Ano63d, Ano63e, Ano64b, Ano64c, Ano64d, Ano64e, Ano65d, Ano65e, Ano65f, Ano65g, Ano65h, Ano66c, Ano66d, Ano66e, Ano66f, Ano66g, Ano67c, Ano67d, Ano67e, Ano67f, Ano67g, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f]. Base [Aub61]. Based [Bra67, Lit67]. Basic [Sch63b, Was63]. Bathtub [Bak67]. Bayes [Har63]. Bayesian [Ans61]. Be [Pet65]. Bean [Sha67]. Behavior [Fer66, Fer67, Kat67]. Behrens [Els69]. Bershad [GT69]. Between [Bli69, Bon69, Duc69, Edw69, RC69, Sea69, Har65, Nat60]. Bias [FU64, Smi69]. Bidwell [Wor65]. Binomial [BHW68, Hog68]. Biologists [Bli69]. Biomedical [Bah68]. Biometric [Ano64-34, Ano65-45, Ano66-45, Ano69-34]. Biometricians [Zel69]. Biometrics [Ano65-45, Ano66-45, Ano69-34, GMN66]. Biometry [Tay66]. Birth [GS65, Ric67]. Birthday [Geh68, Nau68]. Bivariate [HAB69]. Block [Add69]. Boston [Col65]. Both [Ker66]. British [Cha68]. Bulb [Man69b]. Burdick [Ano61k]. Bureau [Men68]. Burgess [Eck69, Ano69-35]. Bursk [Ham63]. Business [Cox65, Har65, Lee60, Sch63b]. Butterbaugh [Pab60]. Buying [LB69]. By-Laws [Ano62-40, Ano63k]. Byer [Ano65-27].

C [Ano64j, Mar60]. C. [Kru68]. Calculation [Stu63]. Calculus [Dar69]. Canada [Ano61s, Ano62u, Ano62v, Ano62w, Ano63u, Ano64t, Ano65z, Ano66y, Bea60, MWR60]. Cancer [Ber63, NFE68]. Card [Hor60]. Career [Ips65]. Careers [Smi68a]. Case [Bli69]. Cautionary [Mih69]. Celebration [Ano65a]. Census [Ano69o, Cor63a, SBB^{+}65]. Centenary [HB67]. Center [Dun67, GRW67]. Central [AR67, Edg63, Mih69]. Century [Ano65c]. Certain [Kul68]. Challenge [Bro61, Rob66]. Challenges [DG69]. Chapter [Ano60e, Ano60f, Ano60g, Ano60h, Ano60i, Ano61f, Ano61g, Ano61h, Ano61i, Ano61j, Ano62f, Ano62g, Ano62h, Ano62i, Ano62j, Ano63f, Ano63g, Ano63h, Ano63i, Ano63j, Ano64f, Ano64g, Ano64h, Ano64i, Ano65i, Ano65j, Ano65k, Ano65l, Ano65m, Ano66h, Ano66i, Ano66j, Ano66k, Ano66l, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k]. Characteristics [Jaf60, RBB65]. Charles [Kin66]. Chi [AR67]. Chi-Squared [AR67]. Choice [GD68]. Class [MP68, Spr69, Ter68]. Classical [Duc69]. Classifications [Stu63]. Classroom [Kna62, Mos62]. Coats [Ste60b]. Coefficient [Sea67]. Coefficients [Duc69, Li64, RC69]. Colleges [Ban60a]. Column [Ano61-35, Bow64a, Bow64b, Bow63a, Bow63b, Bow63c, Bow63d, Bow63e, Gai61a, Gai61b, Gai61c, Gai61d, Han60a, Han60b, Han60c, Han60d, Han60e, Hau62a, Hau62b, Hau62c, Hau62d, Hau62e, Wal65b, Wal65c]. Combinatorial [Dub65, Wis66]. Comments [Bow64c, CFG63]. Committee [Ano62-29, Ano64-35, Ano69o, HHH65, MS69, SBB^{+}69, TMW67, Wer66, Ano69l, Con65, SBB^{+}65]. Communication [Bli69]. Comparison [BHW68, Klo69, RC69, Sak69]. Comparisons [Bra67]. Complement [Cox62]. Complete [Ban61]. Components [Hai66]. Composition [Ano66-43]. Computation [And61, Ano69m, Lie61]. Computer [Goo68, Joh65, Men68]. Computerized [CC69]. Computers [BGB68, Bla68, Hah68, Van68]. Computing [Ban61, Kur68, Cha68]. Conception [Ric67]. Conditions [Wol65]. Conference [Ano69m]. Confidence [Gue69, McH61, Nat60]. Confidential [Cor63a]. Connor [Ano65-46]. Considerations [Ano62-41]. Constitution [Ano62-40, Ano63k]. Construct [Par69]. Consulting [Har63]. Consumer [Fer66, Fer67, Kat67]. Continental [Mos62]. Continuity [Gri67, MG68]. Contrasts [Klo69]. Controversy [Bro63]. Convexity [Gur68]. Cooperation [MS69, You63]. Coordinate [Mey69]. COPSS [HHH65]. Corrado [Ano65n]. Correction [Ano66m, Ano68l, Bro64a, Cur68a, Gri67, MG68]. Correlation [Ano68n, BP68, Bra63, Duc69, Gib68, IG67, RC69, Sea69]. Corrigenda [Ano67m]. Costa [Sas61]. Council [TMW67]. Counting [Mul69]. Course [Cra61, Mos62, Sch63b, Was63]. Covariance [Ser67]. Craps [Smi68b]. Criteria [Ano69n, DFSM65]. Crops [Bea67a]. Croxton [Mar60]. Crump [Dut63]. Curious [Can69]. Current [DG69]. Curve [WD69]. Cut [Hud68]. Cuts [Whi64].

Dana [WCC^{+}60]. Data [CC69, Dun67, GRW67, IG67, Kna62, SS66]. Data-Dredging [SS66]. December [Ano64u, Ros64]. Decision [Lit67, Shu69, Web65]. Decision-Making [Web65]. Decisions [Har65]. Degree [BH68]. Demand [RF60]. Demographic [Aub61]. Density [Hoy68]. Departments [Ler68]. Dependency [Tur62]. Dependent [Rot67]. Depleted [Nel67]. Derivation [Arm67]. Derived [Gur68]. Deriving [Dar69]. Design [Add69, GRW67, JH67]. Determination [McH61]. Development [Ban60b, Kot65, Mal62]. Deviation [Bol68, Bru69, Cur68a, Cur68b, JR68, Mar68a, Mar68b]. Deviations [Sem64]. Device [Coh69, Wat65]. Dewhurst [Ano67z]. Diagramming [Hul66]. Diagrams [Bra63]. Difference [Bon69]. Differences [Edg63]. Different [Sak69]. Diplomacy [Fre63a]. Discrete [Dub66, San67]. Discussant [CFG63]. Discussion [Wer66]. Distinction [Har65]. Distribution [Ahu68, Ano66o, Ano66-41, Bai69, BHW68, Che68, HAB69, Hog68, Mar68a, Mar68b, Nem69, PD69, Ras69, San67, SH68]. Distribution-Free [Nem69]. Distributions [AR67, Dub66, Kul68]. Dogma [Bro61, Rob66]. Dorn [Cor63b]. Dorothy [Ano65o]. Double [Goo69b]. Double-Headers [Goo69b]. Douglass [Ano61k]. Dr. [Har63, Ken65]. Dredging [SS66]. Dublin [Lew69a]. Durand [WCC^{+}60].

E-11 [Wer66]. E. [Ano61k, WCC^{+}60, Wor68]. Eastern [Ano65-45, Ano66-45, Ano69-34]. Easy [Gar68]. Economic [Ano69p, Ano69-33, Ano69-36, McC69, Men68, Tho67, Zar69]. Economics [Cox62]. Economist [Sam60]. Editor [AFD^{+}67, BHD^{+}65, BRG^{+}68, BCU^{+}64, Dem62, EH63, FNG^{+}67, Gol61, GFG^{+}65, Goo65, GBC69, GKH^{+}68, GKK^{+}69, GFB^{+}66, HDC63, Ham61, HGH61, HPWQ63, HSG^{+}67, KGvdB^{+}64, KDMS61, KJ62, Kym68, LOG^{+}69, MDS^{+}65, MING66, MNC^{+}66, Ost62, PKC67, PH68, PWBR65, Pra62, PGH^{+}69, QGS^{+}67, RGH66, SLG^{+}64, Sch63a, SBSK66, Tay65, Ury63, VKT^{+}69]. Education [CL63, Lee60, Zel69]. Educational [Ano66-42, Kep63, Mei60]. Edward [Ano64j, Ano68m]. Edwin [Wor65]. Effect [Ajg67, Han68]. Effective [War69]. Effectiveness [Ros68]. Efficiency [Gea66]. Eighths [Ano65c]. Electric [Arm67]. Elementary [Nem69]. Emil [Fre67]. Employment [Bur62]. ENAR [Ano64-34, Ano65-45, Ano66-45, Ano69-34]. Engineering [Ano64-34, Ano65-45, Ano66-45, Ano69-34]. Engineers [And66]. English [Kot65]. Engquist [Ken65]. Enumerative [Har65]. Equations [Kar64, Som67]. Ernest [Ken65]. Errare [Bow64c]. Erratum [Ano68n]. Error [Mar68a, Mar68b, Pet65, Stu69]. Errors [Fre63b, Ker66, Smi64a, SG69]. Estimated [Pet65]. Estimates [Hoy69]. Estimating [Hud68, SG69, TE61]. Estimation [Bol68, Bru69, Cur68a, Cur68b, Gea66, GD68, Jae66, Koo68, Mar68a, Mar68b, Nel67, Stu69]. Estimator [Ajg67, Han68]. Estimators [Can69, Sak69]. Euler [Cox68]. European [O’T66a, O’T66b]. Events [Gib68, Gor64, Rot67]. Example [You63]. Examples [Hoy69]. Exclusive [Gib68]. Exercise [Jar68, Koo68]. Expectation [Beh67, Chi66, Fle66, Gur67, MP66, Pil68, SSV67]. Expectations [Gur68]. Expected [Lit67, Tur62]. Experience [Sch68]. Experimental [Bon69]. Extension [Nau68]. Extra [Por67]. Extra-Point [Por67].

F [Col65, Kin66, WL61]. Factor [Arm67]. Faculty [Ano66b, Ano66n]. Family [Ano66o]. Federal [Ano61l, Ano61m, Ano62k, Ano62l, Ano62m, Ano62n, Ano62o, Ano63l, Ano63m, Ano63n, Ano63o, Ano64k, Ano64l, Ano64m, Ano64n, Ano65p, Ano65q, Ano65r, Ano65s, Ano65t, Ano66o, Ano66p, Ano66q, Ano66r, Ano66s, BTGK60, BKH61, KCSP61, LCD60, RK60, Smi63, SBFC60, WGW^{+}61, WDC^{+}60, Ano69n, GRW67, You63]. Feedback [Smi64b]. Fellow [Ano62-41]. Fide [BM64]. Field [DS64]. Fields [Ano66-44]. Fieller [Els69]. Fifteen [MS69]. Fiftieth [Ano69o]. Filtering [Tan60]. Finite [Coh69]. Firm [Smi64c]. First [Ano69p, Cra61, Sha69]. Fisher [Els69]. Fitting [Ker66, WD69]. Football [Por67]. Force [Har62, KHH^{+}69]. Forecast [Smi64b]. Forecasts [Zar69]. Formations [Aub61]. Forms [Hul66]. Formulae [Smi69]. Formulas [Pet65]. Foundations [Bir62, Was63]. Fourfold [Mit69]. Francis [Ric64, Joh68]. Frank [Bra66, SK68]. Fred [Cor63b, Mar60]. Frederic [Ano67z]. Free [Nem69]. Front [Ano60j, Ano60k, Ano60l, Ano60m, Ano60n, Ano61n, Ano61o, Ano61p, Ano61q, Ano61r, Ano62p, Ano62q, Ano62r, Ano62s, Ano62t, Ano63p, Ano63q, Ano63r, Ano63s, Ano63t, Ano64o, Ano64p, Ano64q, Ano64r, Ano64s, Ano65u, Ano65v, Ano65w, Ano65x, Ano65y, Ano66t, Ano66u, Ano66v, Ano66w, Ano66x, Ano67n, Ano67o, Ano67p, Ano67q, Ano67r, Ano68o, Ano68p, Ano68q, Ano68r, Ano68s, Ano69q, Ano69r, Ano69s, Ano69t, Ano69u]. Function [Hoy68, Pil68]. Functional [Man69a]. Functions [Bra67, Kul68]. Further [SSV67]. Future [Gor64].

Game [McC69, Smi68b]. Generalized [Add69]. Generation [CC69]. Geographic [Ano66-43]. Geological [Tan60]. Gestation [Ric67]. Gini [Ano65n]. Gladys [Ano67s]. Goals [Was63]. Government [KHH^{+}69, Sub63, Ano61s, Ano62u, Ano62v, Ano62w, Ano63u, Ano64t, Ano65z, Ano66y, Ano69n, MWR60]. Governmental [You63]. Grade [Ano62-41]. Gram [Wat65]. Grant [Pab60]. Graphic [Lie61]. Graphical [Dub66]. Greece [CL63]. Gumbel [Fre67]. Guterman [Ano67t].

Hagood [Tae63]. Hamilton [Ste60b]. Harold [And60, Cor63b, Mad60, Sam60]. Hawaii [Sch68]. Headers [Goo69b]. Health [Bah68, DG69, Ler68]. Henry [Cox62]. Herman [Ano65-27, Ano67t]. Heterovariance [Gea66]. Hinton [Ano65o]. Historical [Goo68]. History [Jar68]. Hotelling [And60, Mad60, Sam60]. Household [Aub61]. Hurdles [Ips65]. Hypergeometric [BHW68]. Hypothesis [Bon69].

IASI [Ric67]. Idea [Dun67]. Implementation [Sch63b]. Incidence [Goo69b]. Incipit [Bow64c]. Income [Ano66o]. Inconsistency [Rot67]. Increasing [Ajg67, Han68]. Independence [Gib68, PD69]. Independent [Mul67]. Index [Ano64u]. India [BS60, Sub63]. Indicators [Coh68]. Industrial [Smi64c]. Inequalities [Gur68]. Inequality [Ano68l, Gur67, Kir68]. Inference [McH61]. Infinite [Rob69]. Information [SH68]. Institute [Ano64-34, Ano65-45, Ano66-45, Ano69-34, Lun68]. Instruction [Ban60a, Tar67]. Instructional [Joh65]. Instructive [Hoy69]. Insurance [Ano62-29]. Integer [Par69]. Integrity [Fre63a]. Interchangeable [KY67]. Interest [Eis65, Eis67, EB67]. Interface [Goo68]. International [Ano60o, Ano60p, Ano60q, Ano60r, Ano60s, Ano61t, Ano61u, Ano61v, Ano61w, Ano61x, Ano62x, Ano62y, Ano62z, Ano62-27, Ano62-28, Ano63v, Ano63w, Ano63x, Ano63y, Ano63z, Ano64v, Ano64w, Ano64x, Ano64y, Ano65-28, Ano65-29, Ano65-30, Ano65-31, Ano65-32, Ano65-44, Ano66z, Ano66-27, Ano66-28, Ano66-29, Ano66-30, Ano67u, Ano67v, Ano67w, Ano67x, Ano67y, Ano68t, Ano68u, Ano68v, Ano68w, Ano68-34, Ano69v]. Interpolation [Lau65]. Interval [McH61]. Intervals [Gue69, Nat60]. Inventory [Tay66]. Invocation [Mih69]. Irrational [Pil68]. Isbell [Mur62]. Island [Sch68]. Issue [Dun67]. Italy [Bar60].

J [Fre67, Ken65, Mor66, Wor68]. J. [Ham63]. Jacob [Cla68]. James [Ano67z]. January [Ano66n]. Japan [BA60]. Jarman [Tae63]. Joint [Ano62-42, Ano63-37, Ano64a, Ano65b, Ano66a, Ano67a, Ano68a, Ano69a, Edg63]. Joseph [B.63]. Journal [GS65]. July [Ros64].

Kendall [Lie61]. Kenichi [Dem65]. Key [Mah65]. King [Mur62]. Kitchen [Bak66]. Knowledge [GD68, Rub68f]. Known [Pet65, Sea67]. Knudsen [Ano65-33]. Koyanagi [Dem65].

L [Ano67s, Her63]. Labor [Men68, Shu69]. Lags [Ste60a]. Largest [Mey69]. Law [Smi64b]. Laws [Ano62-40, Ano63k]. Layout [Klo69]. Learning [War69]. Least [Dar69, Duc69, Sie68]. Least-Squares [Dar69, Sie68]. Lee [Dut63]. Letter [Dem62, EH63, Gol61, Goo65, HGH61, Ost62, Pra62, Sch63a, Ury63]. Letters [AFD^{+}67, BHD^{+}65, BRG^{+}68, BCU^{+}64, FNG^{+}67, GFG^{+}65, GBC69, GKH^{+}68, GKK^{+}69, GFB^{+}66, HDC63, HPWQ63, HSG^{+}67, KGvdB^{+}64, KDMS61, KJ62, Kym68, LOG^{+}69, MDS^{+}65, MING66, MNC^{+}66, PKC67, PH68, PWBR65, PGH^{+}69, QGS^{+}67, RGH66, SLG^{+}64, SBSK66, Tay65, VKT^{+}69]. Life [Pri67, Wol65]. Light [Kot65, Man69b]. Likelihood [Cor66, Hoy69]. Lila [Ano65-33]. Limit [Mih69]. Limits [Ter68]. Line [Ker66]. Linear [Can69, Li64, WD69]. Linguistic [Bro63]. Living [Ros66]. Lloyd [Ano67-27]. Location [JH67]. Logarithmic [Ahu68]. Lottery [LB69]. Louis [Lew69a]. Lowell [Mor66]. Lung [Ber63].

M. [B.63]. Machine [Hor60, Arm67]. Maintain [Fre63a]. Major [Mos67]. Majors [Hog69]. Making [Shu69, Web65]. Manuscripts [Sch67]. March [Ric64]. Margaret [Tae63]. Martin [Hau68]. Masters [BH68]. Materials [Wer66, oH69]. Mathematical [Ano64-34, Ano65-45, Ano66-45, Ano69-34, Bah68, Cra61, Sam60]. Matrix [SH68]. Matter [Ano60b, Ano60c, Ano60d, Ano60j, Ano60k, Ano60l, Ano60m, Ano60n, Ano61a, Ano61b, Ano61c, Ano61d, Ano61e, Ano61n, Ano61o, Ano61p, Ano61q, Ano61r, Ano62a, Ano62b, Ano62c, Ano62d, Ano62e, Ano62p, Ano62q, Ano62r, Ano62s, Ano62t, Ano63a, Ano63b, Ano63c, Ano63d, Ano63e, Ano63p, Ano63q, Ano63r, Ano63s, Ano63t, Ano64b, Ano64c, Ano64d, Ano64e, Ano64o, Ano64p, Ano64q, Ano64r, Ano64s, Ano65d, Ano65e, Ano65f, Ano65g, Ano65h, Ano65u, Ano65v, Ano65w, Ano65x, Ano65y, Ano66c, Ano66d, Ano66e, Ano66f, Ano66g, Ano66t, Ano66u, Ano66v, Ano66w, Ano66x, Ano67c, Ano67d, Ano67e, Ano67f, Ano67g, Ano67n, Ano67o, Ano67p, Ano67q, Ano67r, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano68o, Ano68p, Ano68q, Ano68r, Ano68s, Ano69b]. Matter [Ano69c, Ano69d, Ano69e, Ano69f, Ano69q, Ano69r, Ano69s, Ano69t, Ano69u]. Max [GT69]. Maximization [Hai69]. Maximum [Hoy69]. Mayor [Col65]. McIntyre [Joh68]. Mean [JR68, Mar68a, Mar68b, Ros68, Stu69].

Mean-Square-Error [Mar68a, Mar68b, Stu69]. Meaning [BM62]. Means [Arm67, Bon69, Sea69]. Measure [IG67, Ros68]. Measurement [BM62]. Medical [Ips65, Sch67]. Meehan [B.63]. Meeting [Ano60a, Ano60-30, Ano61-36, Ano62-42, Ano62-29, Ano63-37, Ano64a, Ano64-34, Ano65b, Ano66a, Ano67a, Ano68a, Ano69a, Ano67b]. Meetings [Ano64a, Ano65-45, Ano66-45, Ano69-34, Ano62-42, Ano63-37, Ano65b, Ano66a, Ano67a, Ano68a, Ano69a]. Members [Ano60y, Ano60z, Ano60-27, Ano60-28, Ano60-29, Ano66-42, Ano66-44, GMN66]. Membership [Ano66-41, Ano66-43, Con65, Jaf60, RBB65]. Meteorology [Lun63]. Method [Hud68, SH68, Som67]. Methodological [Ano66-44]. Methods [HB67, Lun63]. Midpoints [Ter68]. Migration [Sch68]. Minimal [SG69]. Minimizing [Sem64]. Minimum [Can69, JH67, Mar68a, Mar68b, Sha69]. Minor [Jar68]. Missing [Jae66]. Models [War69]. Molina [Ano64j]. Moment [Sha69]. Moments [Bai69]. Monetary [Lit67]. Monotonicity [Gur68]. Monroe [Her63]. Moore [Cox62]. Morgenstern [Bow64c]. Mortimer [Lew69b]. Mr [Bea67b]. Mr. [Sha67]. Multiple [GL68, GD68, Klo69, Kra66, Som67]. Multiple-Choice [GD68]. Multivariate [Ano65-44, Ano68-34, SH68]. Must [Pet65]. Mutually [Gib68].

Nath [Ros64]. National [Sch61, Shu69, Con65, Dun67]. Nature [Wol65]. NBER [Zar69]. Nearly [Nem69]. Negligible [Bon69]. Negro [Had67]. Network [Pri67]. News [Ano60y, Ano60z, Ano60-27, Ano60-28, Ano60-29, Ano60t, Ano60u, Ano60v, Ano60w, Ano60x, Ano61y, Ano61z, Ano61-27, Ano61-28, Ano61-29, Ano61-30, Ano61-31, Ano61-32, Ano61-33, Ano61-34, Ano62-30, Ano62-31, Ano62-32, Ano62-33, Ano62-34, Ano62-35, Ano62-36, Ano62-37, Ano62-38, Ano62-39, Ano63-27, Ano63-28, Ano63-29, Ano63-30, Ano63-31, Ano63-32, Ano63-33, Ano63-34, Ano63-35, Ano63-36, Ano64z, Ano64-27, Ano64-28, Ano64-29, Ano64-30, Ano64-31, Ano64-32, Ano64-33, Ano65-34, Ano65-35, Ano65-36, Ano65-37, Ano65-38, Ano65-39, Ano65-40, Ano65-41, Ano65-42, Ano65-43, Ano66-31, Ano66-32, Ano66-33, Ano66-34, Ano66-35, Ano66-36, Ano66-37, Ano66-38, Ano66-39, Ano66-40, Ano67-28, Ano67-29, Ano67-30, Ano67-31, Ano67-32, Ano67-33, Ano67-34, Ano67-35, Ano67-36, Ano67-37, Ano68x, Ano68y, Ano68z, Ano68-27, Ano68-28, Ano68-29, Ano68-30, Ano68-31, Ano68-32, Ano68-33, Ano69w]. News [Ano69x, Ano69y, Ano69z, Ano69-27, Ano69-28, Ano69-29, Ano69-30, Ano69-31, Ano69-32]. Next [Wal65a]. No [Gar68]. Nominal [IG67]. Non [AR67, Lie61, Pet65, WD69]. Non-Central [AR67]. Non-Graphic [Lie61]. Non-Linear [WD69]. Non-Occurrences [Pet65]. Noncentral [Bai69]. Norden [Her63]. Normal [Che68, Hoy68, Mar68a, Mar68b, PD69, SH68, Stu69]. Notation [HHH65, Kra66, HHH65]. Note [Ano68n, Aub61, BP68, Bru69, Fle66, Gea66, Geh68, Hai66, Hai69, Ham61, HAB69, Har65, Ifr69, Jaf60, Joh65, Kar64, Kru68, Mih69, Mul67, Mul69, MP66, NFE68, Pil68, RC69, Ras69, San67, Sea67, Ser67, Spr69, WW69, WD69]. Notes [Ano60e, Ano60f, Ano60g, Ano60h, Ano60i, Ano61f, Ano61g, Ano61h, Ano61i, Ano61j, Ano62f, Ano62g, Ano62h, Ano62i, Ano62j, Ano63f, Ano63g, Ano63h, Ano63i, Ano63j, Ano64f, Ano64g, Ano64h, Ano64i, Ano65i, Ano65j, Ano65k, Ano65l, Ano65m, Ano66h, Ano66i, Ano66j, Ano66k, Ano66l, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k]. Notice [Ano68n]. numbers [Goo69a]. Numerare [Bow64c].

O [Wat65]. Observation [Ker66]. Observations [Hog69]. Obtaining [SH68]. Occasions [Coh69]. Occupancy [Git69, MP68, Spr69]. Occurrences [Pet65]. October [Ric64]. Off [Smi64a]. Office [LS61]. Official [DS64]. One [Hud68, Klo69, Rob66]. One-Way [Klo69]. Only [Hud68]. Operations [Bak66, Bak67]. Opportunities [BGB68, Bla68]. Order [Bra63, Bra67, Hud68, JH67, Kna62]. Ordered [Mey69, Stu63]. Organization [Ban60a]. Organizational [Ban61]. Original [RC69]. Oskar [Bow64c]. Other [Ano69n, Bak66, Bak67]. Outlook [Ano69p, Ano69-33, Ano69-36].

P. [Tae64]. Pairing [Had67]. Palmer [Ano67s]. Paper [Han68, Smi64a]. Paradox [Fre65]. Parameters [Hud68]. Parker [Ham63]. Particle [Mul69]. Parts [Ano68n, BP68, Sea69]. Party [Cha68]. Past [Gor64]. Pedagogical [Wat65]. Percentages [NFE68]. Perimeters [Rub68f]. Perlman [Cla68]. Permutation [JF67]. Personal [Ano61-30, Ano61-31, Ano61-32, Ano61-33, Ano61-34, Ano62-35, Ano62-36, Ano62-37, Ano62-38, Ano62-39, Ano63-32, Ano63-33, Ano63-34, Ano63-35, Ano63-36, Ano64-30, Ano64-31, Ano64-32, Ano64-33, Ano65-39, Ano65-40, Ano65-41, Ano65-42, Ano65-43, Ano66-36, Ano66-37, Ano66-38, Ano66-39, Ano66-40, Ano67-33, Ano67-34, Ano67-35, Ano67-36, Ano67-37, Ano68-29, Ano68-30, Ano68-31, Ano68-32, Ano68-33, Ano69-28, Ano69-29, Ano69-30, Ano69-31, Ano69-32, Dun67]. Personnel [Ler68]. Perspective [Goo68]. Philippines [Ban60b]. Physical [Ano64-34, Ano65-45, Ano66-45, Ano69-34]. Pillai [Kru68]. Plan [McC69]. Plans [Ban61]. Play [FW64]. Plotting [Bra63]. Point [Hog68, Por67, Tar67]. Points [Bra63]. Poisson [HAB69]. Policy [Coh68, Kep63, McC69]. Polynomial [JH67]. Pools [Wil68]. Population [Hog68, Nel67]. Populations [Coh69]. Positive [SSV67]. Post [LS61]. Practical [Ban61]. Precision [Ajg67, Han68]. Predicted [WW69]. Preface [Rub60a]. Preparation [Bah68]. Present [Ker66, Sub63]. Presentational [Wat65]. Preservation [oH69]. President [Ano61-35, Bow64a, Bow64b, Bow63a, Bow63b, Bow63c, Bow63d, Bow63e, Gai61a, Gai61b, Gai61c, Gai61d, Han60a, Han60b, Han60c, Han60d, Han60e, Hau62a, Hau62b, Hau62c, Hau62d, Hau62e, Mos67, Wal65b, Wal65c]. Presidents [Ano69l]. Price [LB69]. Prime [Pri67]. Principal [Hai66]. Principle [Cor66, Rot67]. Principles [DFSM65]. Priority [Cur68a, Mar68b]. Privacy [Dun67]. Probability [Ano64-35, GS65, Gor64, Gri68, Hoy68, Mos62, O’T66a, O’T66b, Smi68b, Tho67]. Problem [Dub65, Els69, Geh68, Git69, Nau68, Rob69, War69, Wis66]. Problems [DS64, MP68, Smi68b, Spr69]. Procedures [SS66]. Prochnow [Ano67-27]. Professor [Tar67]. Profile [Ano66-41, Ano66-42, Ano66-43, Ano66-44, HB67]. Profile-Age [Ano66-41]. Profile-Educational [Ano66-42]. Profile-Geographic [Ano66-43]. Profile-Statistical [Ano66-44]. Profit [Rot67]. Program [Sch67, Ano60-30, Ano61-36, Ano64-34, Ano65-44, Ano68-34, Har62]. Programmed [Tar67]. Programming [Par69]. Programs [Hog69, Pri67]. Proof [Klo69, Nem69, SG69]. Property [KY67, Sha69]. Proposed [Ano62-40]. Public [Bah68, Coh68]. Publications [SBB^{+}69]. Published [Ano66o]. Pure [Cox62]. Puzzle [Fre65].

Quadratic [Hul66]. Quarter [Ano69p, Ano69-33, Ano69-36]. Questions [Ano66m, Ano67m, Rub60b, Rub60c, RS60, Rub60d, Rub61, RE61, RR61, RH61, RS61, Rub62a, Rub62b, Rub62c, Rub62d, Rub62e, Rub63a, Rub63b, RPG^{+}63, RCF63, Rub63c, Rub64a, Rub64b, Rub64c, Rub64d, Rub65a, Rub65b, RR65, Rub65c, Rub65d, RH66, RA66, RNWR66, Rub66a, Rub66b, RNCE67, Rub67a, RR67, Rub67b, RM67, Rub68a, Rub68b, Rub68c, Rub68d, Rub68e, Rub69a, Rub69b, Rub69c, RG69, Rub69d]. Qui [Bow64c].

Randolph [Ano65-33]. Random [Beh67, Chi66, Fle66, Gur67, Gur68, Had67, Ker61, KY67, Mul67, MP66, Pil68, SSV67, Ser67, Goo69a]. Randomized [Add69]. Range [JR68]. Rank [Bra63, Kna62]. Rank-Order [Kna62]. Ratio [Koo68, Mul67]. Ratios [Smi69]. Really [Bro64a, Bro64b]. Recession [Moo67]. Reciprocal [Chi66, Fle66, Gur67, MP66, SSV67, Ser67]. Recommended [HHH65]. Recruitment [DS64]. Reduced [Stu69]. Reed [Mor66]. Reflections [Bur62]. Regarding [Mih69]. Region [Ano65-45, Ano66-45, Ano69-34]. Regional [Ano64-34, Ano65-45, Ano66-45, Ano69-34, MS69]. Regression [Duc69, Fre63b, Gea66, GL68, Hai66, Kar64, Kra66, Li64, Smi64a, Som67, WW69, War69]. Rejoinder [Bea67b, Fer67]. Relation [Edw69, Nat60]. Relationship [Duc69]. Relative [Hul66]. Remarks [Col65, Han68, Ker61, SSV67, Smi64a]. Report [And66, Ano64-35, BM62, Con65, Cox65, DFSM65, KHH^{+}69, SBB^{+}65, SBB^{+}69, Tay66, Wer66]. Reports [Cor63a, Mos67]. Representatives [Ano66b, Ano66n]. Required [Sch63b, TE61, Was63]. Requirements [BH68]. Research [And60, Bak66, Bak67, Har62, Kur68, Mal62, TMW67]. Residual [Gea66]. Response [NFE68, Rob66, WW69]. Response-to-Treatment [NFE68]. Results [Can69]. Reviewing [Sch67]. Revolution [Bea67a]. Rho [Stu63]. Rica [Sas61]. Rice [Ano69-35, Sib69]. Rice/Robert [Ano69-35]. Robert [Ano69-35, Eck69, Ste60b]. Role [Kur68]. Room [Har63]. Roommate [Had67]. Root [MP66]. Ross [SK68]. Round [Smi64a, Wer66]. Round-Off [Smi64a]. Roundoff [Fre63b]. Roy [Ros64]. Russian [Kot65].

S [Ano65-46, Kru68, Mos64]. S. [Dut63]. Samarendra [Ros64]. Sample [Ajg67, Bra67, Han68, Hog68, McH61, TE61]. Samples [Stu69]. Sampling [Ano68l, Coh69, Kir68, Nel67, Tan60]. Samuel [Hau60, Mos64]. Sarle [Kin66]. Satisfied [Gur67]. Scatter [Bra63]. Scheffé [Els69, Klo69]. School [Ano64-35]. Schools [Cox65, O’T66a]. Science [Bah68, Goo68, TMW67, Van68]. Sciences [Ano64-34, Ano65-45, Ano66-45, Ano69-34]. Second [Ano68-34, Ano69-33]. Secondary [Ano64-35, O’T66a]. Section [Ano64-34, Ano65-45, Ano66-45, Ano69-34, GMN66]. Sed [BM64]. Selling [LB69]. Sequential [Cor66, Ifr69]. Series [Ahu68, Wil68]. Service [Ban61]. Set [Hud68]. Sex [Gri68]. Sharing [Hah68, Kur68]. Shewhart [Dem67]. Short [Hud68]. Shortest [Gue69]. Significance [GL68, Nat60]. Significant [Rot67]. Simple [Duc69, Hoy68, Klo69, Kra66, SH68, Som67]. Simplified [Som67]. Simulation [Sak69]. Sink [Bak66]. Size [Ajg67, Bra67, Han68, McH61, TE61]. Smoking [Ber63]. Smoothing [Ste60a]. Social [Ano62-29, Coh68, Van68, TMW67]. Societies [Ano69l]. Society [Ano64-34, Ano65-45, Ano66-45, Ano69-34]. Solution [Els69, Wis66]. Solutions [Dub65]. Solving [Som67]. Some [Ano62-41, Bur62, Can69, Che68, Hah68, Han68, Hog69, O’T66b, Pet65, Smi68b]. Southern [MS69]. Spain [Poc60]. Spearman [Stu63]. Special [BM62]. Spiegelman [Lew69b]. Spring [Ano65-45, Ano66-45, Ano69-34]. Square [Mar68a, Mar68b, MP66, Stu69]. Squared [AR67, SG69]. Squares [Dar69, Duc69, Sie68]. Stamp [Rob69]. Standard [Bol68, Bru69, Cur68a, Cur68b, Hoy68, Mar68a, Mar68b]. Standards [HHH65, Web65]. State [Ler68, Sch68, Sub63, You63]. Statehood [Sch61]. Statesman [Mos64]. Statist [WL61]. Statistic [Ras69]. Statistical [Ano60o, Ano60p, Ano60q, Ano60r, Ano60s, Ano60-30, Ano61l, Ano61m, Ano61s, Ano61t, Ano61u, Ano61v, Ano61w, Ano61x, Ano61-36, Ano62-42, Ano62k, Ano62l, Ano62m, Ano62n, Ano62o, Ano62u, Ano62v, Ano62w, Ano62x, Ano62y, Ano62z, Ano62-27, Ano62-28, Ano62-40, Ano63-37, Ano63k, Ano63l, Ano63m, Ano63n, Ano63o, Ano63u, Ano63v, Ano63w, Ano63x, Ano63y, Ano63z, Ano64a, Ano64k, Ano64l, Ano64m, Ano64n, Ano64t, Ano64v, Ano64w, Ano64x, Ano64y, Ano65b, Ano65p, Ano65q, Ano65r, Ano65s, Ano65t, Ano65z, Ano65-28, Ano65-29, Ano65-30, Ano65-31, Ano65-32, Ano65-45, Ano66a, Ano66p, Ano66q, Ano66r, Ano66s, Ano66y, Ano66z, Ano66-27, Ano66-28, Ano66-29, Ano66-30, Ano66-44, Ano66-45, Ano67a, Ano67u, Ano67v, Ano67w, Ano67x, Ano67y, Ano68a]. Statistical [Ano68t, Ano68u, Ano68v, Ano68w, Ano69a, Ano69l, Ano69m, Ano69v, Ano69-34, Ban60a, Ban60b, BTGK60, BKH61, Bro61, Bro63, Cha68, CL63, Cox68, DS64, Dub65, FW64, GRW67, Hah68, HHH65, Ips65, Jaf60, KCSP61, Kot65, Kur68, LCD60, Ler68, Lun63, MWR60, Par69, RK60, Rob66, RBB65, Sch67, SBB^{+}65, Smi63, Smi64c, SBFC60, TMW67, Wal65a, Web65, Wis66, WGW^{+}61, WDC^{+}60, Cox62]. Statistician [Ano64u]. Statisticians [Bah68, BGB68, Bla68, Bli69, DG69, Eis65, Eis67, EB67, Fre63a, RF60, Van68, Whi64, Zar69]. Statistics [And60, And66, Ano62-29, Ano64-34, Ano64-35, Ano65-45, Ano66o, Ano66-45, Ano69-34, Ans61, BA60, Bar60, Bea60, Bir62, BM62, Bra67, BS60, CC69, Coh68, Cox65, Cra61, Duc69, Fer66, Fer67, Goo68, Har62, Hof61, Hog69, Hud68, Joh65, Kat67, KHH^{+}69, Kep63, Kot65, Lee60, LS61, Lun68, Mah65, Mal62, Man69b, Mei60, Mos62, Mos64, O’T66a, O’T66b, Poc60, R.68, Rub68f, Sas61, Sch63b, Sch61, Sch68, Shu69, Smi68a, Sub63, Tar67, Tuk65, Was63, Wer66, Wil65, You63, BH68, Men68, MS69]. Status [Cor63a, O’T66b]. Storage [KHH^{+}69]. Stouffer [Hau60]. Strategy [Por67]. Stuart [Ano69-35, Sib69]. Students [Had67, Tur62]. Studies [Sak69]. Study [Bli69, R.68]. Subcommittee [DFSM65]. Successive [Coh69]. Sum [SG69]. Summary [And66, Cox65]. Supermatrix [Sie68]. Supply [RF60]. Survey [Ano69p, Ano69-33, Ano69-36, SS66, Zar69]. Surveys [Ano69n, Har65, HB67]. Swept [Goo69b]. Swift [Arm67]. Symbols [HHH65, Ros66]. Symmetric [KY67, Kul68]. Symposium [Ano65-44, Ano68-34]. Synthesizing [War69].

Table [Mit69, Pri67, Wer66]. Tables [Gri67, Kul68, Lau65]. Tabulation [Hor60]. Taitel [Hau68]. Task [KHH^{+}69]. Tau [Lie61]. Teacher [Mad60]. Teaching [And66, BA60, Bar60, Bea60, CC69, Cox65, Hof61, Nat60, O’T66a, Poc60, Sas61]. Technical [Tuk65]. Technique [Par69]. Techniques [Ano66-44]. Technology [Mah65]. Television [Pri67]. Ten [Ano65c]. Ten-Eighths [Ano65c]. Tendency [Edg63]. Terminology [Kot65]. Test [Bon69, Edw69, Gri67, Nem69, Ras69, Wat65]. Test-O-Gram [Wat65]. Testing [Mal62, Wer66]. Tests [Dub66, GL68, GD68, Ifr69, Nat60, Par69]. th [JH67, Mey69]. Their [Fre63a]. Theorem [Els69, JF67, Klo69, Mih69, Cox68]. Theorems [Hul66]. Theory [Arm67]. Thesis [BH68]. Third [Ano69-36]. Three [Mos67]. Time [Hah68, Kur68, Pri67, Ros68]. Time-Sharing [Hah68, Kur68]. Tom [Arm67]. Tool [Joh65]. Tools [Tuk65]. Training [Bah68, Fre63a, Ler68, Lun68, Smi64c, Ano62-29, TMW67]. Transformed [RC69]. Treatment [Kar64, NFE68]. Trials [Cor66]. Tribute [Ken65]. Tuple [Mey69]. TV [Mei60, Mos62]. Two [Hog68, Hoy69, Hud68, Li64, Mul67, Stu63]. Two-Point [Hog68]. Two-Way [Stu63].

U.S.S.R. [Kot65]. Unbiased [Bol68, Bru69, Can69, Cur68a, Cur68b]. Undergraduate [Hog69]. Undertaking [Ano69n]. Unemployment [BM62]. Uniform [San67]. Universities [Ban60a, BH68, O’T66b]. University [Ban61]. Unkind [Whi64]. Use [Kna62, Sea67]. Useful [Che68]. Uses [Hah68]. Using [Can69, Els69].

Value [Beh67, Jae66]. Values [Lit67]. Variability [Edg63]. Variable [Chi66, Fle66, Gur67, Kar64, Man69a, MP66, Pil68, SSV67, Ser67]. Variables [Ano68l, Gur68, JH67, Ker66, Kir68, KY67, Mul67, RC69]. Variance [And61, Can69, Hog68, San67, Smi69, WW69]. Variances [Nem69, TE61]. Variation [Sea67]. Vector [Beh67]. Via [SG69, Mei60]. Victories [Gar68]. Vide [BM64]. View [Bir62, Tar67]. Views [Li64]. Voting [FU64]. vs [Gor64, Kot65].

W [Ano69-35]. Waiting [Ros68]. Walter [Dem67, Ric64, WL61]. Warning [Fre63b]. Way [Klo69, Stu63]. Weather [Bea67a, Bea67b, Sha67]. Weights [Dar69]. Western [O’T66a]. Whelpton [Tae64]. White [Ano68m, Had67]. Wholes [Ano68n, BP68, Sea69]. Wilbur [Eck69]. Wilcoxon [Bra66]. Wilks [Mos64]. Willcox [Ric64, WL61]. Willford [Mur62]. William [Ano65-46]. Wilson [Wor65]. within [Smi64c]. without [Bra63, Dar69]. Word [FW64]. Working [Cha68, Wol65]. World [Wil68]. Wrong [Bro64a, Bro64b].

Years [MS69, Wal65a].

Zero [Gib68].

References

Addelman:1969:GRB

[Add69] Sidney Addelman. The generalized randomized block design. The American Statistician, 23(4):35–36, October 1969. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

            http://www.jstor.org/stable/2681737.

Aidenoff:1967:LE

[AFD^{+}67] Layle Silbert Aidenoff, Joseph L. Fleiss, James M. Dickey, Wolfgang Buhler, William Kruskal, Douglas K. Spiegel, William J. Blot, V. R. Rao Uppuluri, Paul W. Eaton, Lawrence J. Remington, Kenneth E. Schnelle, L. R. Shenton, Hans Ury, Robert Brandwein, C. J. Thomas, Roy W. Meadows, Hartwig Harders, and F. L. Abraham. Letters to the Editor. The American Statistician, 21(4):49–55, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

            org/stable/2682109.

Ahuja:1968:LSD

[Ahu68] J. C. Ahuja. On the logarithmic series distribution. The American Statistician, 22(4):37, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

            www.jstor.org/stable/2682023.

Ajgaonkar:1967:EIS

[Ajg67] S. G. Prabhu Ajgaonkar. The effect of increasing sample size on the precision of an estimator. The American Statistician, 21 (4):26–28, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/ stable/2682102. See remarks [Han68].

Anderson:1960:HHR

[And60] T. W. Anderson. Harold Hotelling’s research in statistics. The American Statistician, 14(3):17–21, June 1960. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682167.

Anderson:1961:CAV

[And61] Kenneth E. Anderson. Computation of t by analysis of variance. The American Statistician, 15(2):18–19, April 1961. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682728.

Anderson:1966:TSE

[And66] Virgil L. Anderson. The teaching of statistics for engineers — summary report. The American Statistician, 20(1):14–17, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681670.

Anonymous:1960:AMA

[Ano60a] Anonymous. 1960 Annual Meeting — August 23–26. The American Statistician, 14(3):26–27, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682169.

Anonymous:1960:BMa

[Ano60b] Anonymous. Back matter. The American Statistician, 14(2):33, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682251.

Anonymous:1960:BMb

[Ano60c] Anonymous. Back matter. The American Statistician, 14(3):46, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682174.

Anonymous:1960:BMc

[Ano60d] Anonymous. Back matter. The American Statistician, 14(4): 42, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681386.

Anonymous:1960:CNa

[Ano60e] Anonymous. Chapter notes. The American Statistician, 14(1):35– 37, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682134.

Anonymous:1960:CNb

[Ano60f] Anonymous. Chapter notes. The American Statistician, 14(2): 30–32, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682250.

Anonymous:1960:CNc

[Ano60g] Anonymous. Chapter notes. The American Statistician, 14(3): 41–43, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682172.

Anonymous:1960:CNd

[Ano60h] Anonymous. Chapter notes. The American Statistician, 14(4): 37–39, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681385.

Anonymous:1960:CNe

[Ano60i] Anonymous. Chapter notes. The American Statistician, 14(5):22– 24, December 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682481.

Anonymous:1960:FMa

[Ano60j] Anonymous. Front matter. The American Statistician, 14(1):26– 32, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682122.

Anonymous:1960:FMb

[Ano60k] Anonymous. Front matter. The American Statistician, 14(2): 7–26, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682237.

Anonymous:1960:FMc

[Ano60l] Anonymous. Front matter. The American Statistician, 14(3): 8–44, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682160.

Anonymous:1960:FMd

[Ano60m] Anonymous. Front matter. The American Statistician, 14(4):18– 31, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681373.

Anonymous:1960:FMe

[Ano60n] Anonymous. Front matter. The American Statistician, 14(5): 9, December 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682471.

Anonymous:1960:ISAa

[Ano60o] Anonymous. International statistical activities. The American Statistician, 14(1):8–9, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682125.

Anonymous:1960:ISAb

[Ano60p] Anonymous. International statistical activities. The American Statistician, 14(2):6, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682240.

Anonymous:1960:ISAc

[Ano60q] Anonymous. International statistical activities. The American Statistician, 14(3):9–10, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682163.

Anonymous:1960:ISAd

[Ano60r] Anonymous. International statistical activities. The American Statistician, 14(4):8–10, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681376.

Anonymous:1960:ISAe

[Ano60s] Anonymous. International statistical activities. The American Statistician, 14(5):6–8, December 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682474.

Anonymous:1960:Na

[Ano60t] Anonymous. News. The American Statistician, 14(1):2–7, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682124.

Anonymous:1960:Nb

[Ano60u] Anonymous. News. The American Statistician, 14(2):2–5, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682239.

Anonymous:1960:Nc

[Ano60v] Anonymous. News. The American Statistician, 14(3):2–7, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682162.

Anonymous:1960:Nd

[Ano60w] Anonymous. News. The American Statistician, 14(4):2–40, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681375.

Anonymous:1960:Ne

[Ano60x] Anonymous. News. The American Statistician, 14(5):2–5, December 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682473.

Anonymous:1960:NAMa

[Ano60y] Anonymous. News about members. The American Statistician, 14(1):33–34, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682133.

Anonymous:1960:NAMb

[Ano60z] Anonymous. News about members. The American Statistician, 14(2):27, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682247.

Anonymous:1960:NAMc

[Ano60-27] Anonymous. News about members. The American Statistician, 14(3):40, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682171.

Anonymous:1960:NAMd

[Ano60-28] Anonymous. News about members. The American Statistician, 14(4):33–35, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681383.

Anonymous:1960:NAMe

[Ano60-29] Anonymous. News about members. The American Statistician, 14(5):25, December 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682483.

Anonymous:1960:PAM

[Ano60-30] Anonymous. Program of the 120th Annual Meeting of the American Statistical Association. The American Statistician, 14(3): 28–38, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682170.

Anonymous:1961:BMa

[Ano61a] Anonymous. Back matter. The American Statistician, 15(1): ??, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682511.

Anonymous:1961:BMb

[Ano61b] Anonymous. Back matter. The American Statistician, 15(2): 31–32, April 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682732.

Anonymous:1961:BMc

[Ano61c] Anonymous. Back matter. The American Statistician, 15(3):32, June 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682631.

Anonymous:1961:BMd

[Ano61d] Anonymous. Back matter. The American Statistician, 15(4):43– 44, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681828.

Anonymous:1961:BMe

[Ano61e] Anonymous. Back matter. The American Statistician, 15(5):??, December 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681617.

Anonymous:1961:CNa

[Ano61f] Anonymous. Chapter notes. The American Statistician, 15(1):32– 34, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682509.

Anonymous:1961:CNb

[Ano61g] Anonymous. Chapter notes. The American Statistician, 15(2): 29–30, April 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682731.

Anonymous:1961:CNc

[Ano61h] Anonymous. Chapter notes. The American Statistician, 15(3): 28–30, June 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682630.

Anonymous:1961:CNd

[Ano61i] Anonymous. Chapter notes. The American Statistician, 15(4): 35–37, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681827.

Anonymous:1961:CNe

[Ano61j] Anonymous. Chapter notes. The American Statistician, 15(5):22– 23, December 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681616.

Anonymous:1961:DB

[Ano61k] Anonymous. E. Douglass Burdick, 1905–1961. The American Statistician, 15(1):26, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682506.

Anonymous:1961:FSAa

[Ano61l] Anonymous. Federal statistical activities. The American Statistician, 15(4):9–11, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681820.

Anonymous:1961:FSAb

[Ano61m] Anonymous. Federal statistical activities. The American Statistician, 15(5):8–9, December 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681611.

Anonymous:1961:FMa

[Ano61n] Anonymous. Front matter. The American Statistician, 15(1): 31, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682496.

Anonymous:1961:FMb

[Ano61o] Anonymous. Front matter. The American Statistician, 15(2):??, April 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682721.

Anonymous:1961:FMc

[Ano61p] Anonymous. Front matter. The American Statistician, 15(3):??, June 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682619.

Anonymous:1961:FMd

[Ano61q] Anonymous. Front matter. The American Statistician, 15(4):13– 19, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681815.

Anonymous:1961:FMe

[Ano61r] Anonymous. Front matter. The American Statistician, 15(5): ??, December 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681606.

Anonymous:1961:GCS

[Ano61s] Anonymous. Government of Canada statistical activities. The American Statistician, 15(4):12, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681821.

Anonymous:1961:ISAa

[Ano61t] Anonymous. International statistical activities. The American Statistician, 15(1):8–9, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682501.

Anonymous:1961:ISAb

[Ano61u] Anonymous. International statistical activities. The American Statistician, 15(2):8–10, April 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682725.

Anonymous:1961:ISAc

[Ano61v] Anonymous. International statistical activities. The American Statistician, 15(3):8–10, June 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682622.

Anonymous:1961:ISAd

[Ano61w] Anonymous. International statistical activities. The American Statistician, 15(4):7–8, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681818.

Anonymous:1961:ISAe

[Ano61x] Anonymous. International statistical activities. The American Statistician, 15(5):6–7, December 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681609.

Anonymous:1961:Na

[Ano61y] Anonymous. News. The American Statistician, 15(1):2–7, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682499.

Anonymous:1961:Nb

[Ano61z] Anonymous. News. The American Statistician, 15(2):2–8, April 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682723.

Anonymous:1961:Nc

[Ano61-27] Anonymous. News. The American Statistician, 15(3):2–8, June 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682621.

Anonymous:1961:Nd

[Ano61-28] Anonymous. News. The American Statistician, 15(4):2–6, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681817.

Anonymous:1961:Ne

[Ano61-29] Anonymous. News. The American Statistician, 15(5):2–6, December 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681608.

Anonymous:1961:PNa

[Ano61-30] Anonymous. Personal news. The American Statistician, 15(1): 7–35, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682500.

Anonymous:1961:PNb

[Ano61-31] Anonymous. Personal news. The American Statistician, 15(2):8, April 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682724.

Anonymous:1961:PNc

[Ano61-32] Anonymous. Personal news. The American Statistician, 15(3): 13–31, June 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682624.

Anonymous:1961:PNd

[Ano61-33] Anonymous. Personal news. The American Statistician, 15(4):8– 42, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681819.

Anonymous:1961:PNe

[Ano61-34] Anonymous. Personal news. The American Statistician, 15(5):7– 25, December 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681610.

Anonymous:1961:PC

[Ano61-35] Anonymous. The President’s column. The American Statistician, 15(1):1–20, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682497.

Anonymous:1961:PAM

[Ano61-36] Anonymous. Program of the 121st Annual Meeting of the American Statistical Association. The American Statistician, 15(4):24– 31, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681825.

Anonymous:1962:BMa

[Ano62a] Anonymous. Back matter. The American Statistician, 16(1): ??, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681951.

Anonymous:1962:BMb

[Ano62b] Anonymous. Back matter. The American Statistician, 16(2):??, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681541.

Anonymous:1962:BMc

[Ano62c] Anonymous. Back matter. The American Statistician, 16(3): 47–48, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682077.

Anonymous:1962:BMd

[Ano62d] Anonymous. Back matter. The American Statistician, 16(4): ??, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681438.

Anonymous:1962:BMe

[Ano62e] Anonymous. Back matter. The American Statistician, 16(5):??, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681965.

Anonymous:1962:CNa

[Ano62f] Anonymous. Chapter notes. The American Statistician, 16(1):27– 29, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681950.

Anonymous:1962:CNb

[Ano62g] Anonymous. Chapter notes. The American Statistician, 16(2): 18–21, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681539.

Anonymous:1962:CNc

[Ano62h] Anonymous. Chapter notes. The American Statistician, 16(3): 42–44, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682076.

Anonymous:1962:CNd

[Ano62i] Anonymous. Chapter notes. The American Statistician, 16(4): 35–36, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681437.

Anonymous:1962:CNe

[Ano62j] Anonymous. Chapter notes. The American Statistician, 16(5):31– 33, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681963.

Anonymous:1962:FSAa

[Ano62k] Anonymous. Federal statistical activities. The American Statistician, 16(1):13–15, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681945.

Anonymous:1962:FSAb

[Ano62l] Anonymous. Federal statistical activities. The American Statistician, 16(2):8–9, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681535.

Anonymous:1962:FSAc

[Ano62m] Anonymous. Federal statistical activities. The American Statistician, 16(3):11–45, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682068.

Anonymous:1962:FSAd

[Ano62n] Anonymous. Federal statistical activities. The American Statistician, 16(4):12–13, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681431.

Anonymous:1962:FSAe

[Ano62o] Anonymous. Federal statistical activities. The American Statistician, 16(5):7–8, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681958.

Anonymous:1962:FMa

[Ano62p] Anonymous. Front matter. The American Statistician, 16(1):7– 26, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681940.

Anonymous:1962:FMb

[Ano62q] Anonymous. Front matter. The American Statistician, 16(2):??, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681530.

Anonymous:1962:FMc

[Ano62r] Anonymous. Front matter. The American Statistician, 16(3):18, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682063.

Anonymous:1962:FMd

[Ano62s] Anonymous. Front matter. The American Statistician, 16(4): 31, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681426.

Anonymous:1962:FMe

[Ano62t] Anonymous. Front matter. The American Statistician, 16(5):9– 26, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681952.

Anonymous:1962:GCSa

[Ano62u] Anonymous. Government of Canada statistical activities. The American Statistician, 16(1):16, February 1962. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681946.

Anonymous:1962:GCSb

[Ano62v] Anonymous. Government of Canada statistical activities. The American Statistician, 16(4):11–12, October 1962. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681430.

Anonymous:1962:GCSc

[Ano62w] Anonymous. Government of Canada statistical activities. The American Statistician, 16(5):6–7, December 1962. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681957.

Anonymous:1962:ISAa

[Ano62x] Anonymous. International statistical activities. The American Statistician, 16(1):11–12, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681944.

Anonymous:1962:ISAb

[Ano62y] Anonymous. International statistical activities. The American Statistician, 16(2):7–8, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681534.

Anonymous:1962:ISAc

[Ano62z] Anonymous. International statistical activities. The American Statistician, 16(3):10–11, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682067.

Anonymous:1962:ISAd

[Ano62-27] Anonymous. International statistical activities. The American Statistician, 16(4):10–11, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681429.

Anonymous:1962:ISAe

[Ano62-28] Anonymous. International statistical activities. The American Statistician, 16(5):6, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681956.

Anonymous:1962:MAC

[Ano62-29] Anonymous. Meeting of ASA Committee on Social Insurance Statistics and Training. The American Statistician, 16(3):40–41, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682075.

Anonymous:1962:Na

[Ano62-30] Anonymous. News. The American Statistician, 16(1):2–10, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681942.

Anonymous:1962:Nb

[Ano62-31] Anonymous. News. The American Statistician, 16(2):2–6, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681532.

Anonymous:1962:Nc

[Ano62-32] Anonymous. News. The American Statistician, 16(3):2–9, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682065.

Anonymous:1962:Nd

[Ano62-33] Anonymous. News. The American Statistician, 16(4):2–9, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681428.

Anonymous:1962:Ne

[Ano62-34] Anonymous. News. The American Statistician, 16(5):2–5, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681954.

Anonymous:1962:PNa

[Ano62-35] Anonymous. Personal news. The American Statistician, 16(1):10– 31, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681943.

Anonymous:1962:PNb

[Ano62-36] Anonymous. Personal news. The American Statistician, 16(2): 6–22, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681533.

Anonymous:1962:PNd

[Ano62-37] Anonymous. Personal news. The American Statistician, 16(3): 9–45, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682066.

Anonymous:1962:PNe

[Ano62-38] Anonymous. Personal news. The American Statistician, 16(4): 32–34, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681436.

Anonymous:1962:PNf

[Ano62-39] Anonymous. Personal news. The American Statistician, 16(5):5– 35, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681955.

Anonymous:1962:PNC

[Ano62-40] Anonymous. Proposed new constitution and by-laws of the American Statistical Association. The American Statistician, 16(3): 19–25, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682070.

Anonymous:1962:SCG

[Ano62-41] Anonymous. Some considerations on the grade of Fellow in the ASA. The American Statistician, 16(3):25–28, June 1962. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682071.

Anonymous:1962:AMJ

[Ano62-42] Anonymous. The 122nd Annual Meeting: Joint Statistical Meetings. The American Statistician, 16(3):29–36, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682072.

Anonymous:1963:BMa

[Ano63a] Anonymous. Back matter. The American Statistician, 17(1): ??, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682495.

Anonymous:1963:BMb

[Ano63b] Anonymous. Back matter. The American Statistician, 17(2):??, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685520.

Anonymous:1963:BMc

[Ano63c] Anonymous. Back matter. The American Statistician, 17(3): 55–56, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681351.

Anonymous:1963:BMd

[Ano63d] Anonymous. Back matter. The American Statistician, 17(4):38– 40, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682600.

Anonymous:1963:BMe

[Ano63e] Anonymous. Back matter. The American Statistician, 17(5):28, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685675.

Anonymous:1963:CNa

[Ano63f] Anonymous. Chapter notes. The American Statistician, 17(1):28– 30, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682494.

Anonymous:1963:CNb

[Ano63g] Anonymous. Chapter notes. The American Statistician, 17(2): 31–33, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685518.

Anonymous:1963:CNc

[Ano63h] Anonymous. Chapter notes. The American Statistician, 17(3): 50–52, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681344.

Anonymous:1963:CNd

[Ano63i] Anonymous. Chapter notes. The American Statistician, 17(4): 33–35, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682598.

Anonymous:1963:CNe

[Ano63j] Anonymous. Chapter notes. The American Statistician, 17(5):25– 26, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685674.

Anonymous:1963:CLA

[Ano63k] Anonymous. Constitution and by-laws of the American Statistical Association. The American Statistician, 17(2):25–29, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685516.

Anonymous:1963:FSAa

[Ano63l] Anonymous. Federal statistical activities. The American Statistician, 17(1):13–15, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682489.

Anonymous:1963:FSAb

[Ano63m] Anonymous. Federal statistical activities. The American Statistician, 17(2):10–24, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685511.

Anonymous:1963:FSAc

[Ano63n] Anonymous. Federal statistical activities. The American Statistician, 17(4):12–30, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682592.

Anonymous:1963:FSAd

[Ano63o] Anonymous. Federal statistical activities. The American Statistician, 17(5):8–20, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685668.

Anonymous:1963:FMa

[Ano63p] Anonymous. Front matter. The American Statistician, 17(1):16– 31, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682484.

Anonymous:1963:FMb

[Ano63q] Anonymous. Front matter. The American Statistician, 17(2):34, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685507.

Anonymous:1963:FMc

[Ano63r] Anonymous. Front matter. The American Statistician, 17(3):27, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681330.

Anonymous:1963:FMd

[Ano63s] Anonymous. Front matter. The American Statistician, 17(4):25– 26, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682586.

Anonymous:1963:FMe

[Ano63t] Anonymous. Front matter. The American Statistician, 17(5): 6, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685663.

Anonymous:1963:GCS

[Ano63u] Anonymous. Government of Canada statistical activities. The American Statistician, 17(4):9, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682589.

Anonymous:1963:ISAa

[Ano63v] Anonymous. International statistical activities. The American Statistician, 17(1):10–12, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682487.

Anonymous:1963:ISAb

[Ano63w] Anonymous. International statistical activities. The American Statistician, 17(2):8–9, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685510.

Anonymous:1963:ISAc

[Ano63x] Anonymous. International statistical activities. The American Statistician, 17(3):11–12, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681334.

Anonymous:1963:ISAd

[Ano63y] Anonymous. International statistical activities. The American Statistician, 17(4):10–11, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682590.

Anonymous:1963:ISAe

[Ano63z] Anonymous. International statistical activities. The American Statistician, 17(5):7, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685666.

Anonymous:1963:Na

[Ano63-27] Anonymous. News. The American Statistician, 17(1):2–9, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682486.

Anonymous:1963:Nb

[Ano63-28] Anonymous. News. The American Statistician, 17(2):2–7, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685509.

Anonymous:1963:Nc

[Ano63-29] Anonymous. News. The American Statistician, 17(3):2–10, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681332.

Anonymous:1963:Nd

[Ano63-30] Anonymous. News. The American Statistician, 17(4):2–9, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682588.

Anonymous:1963:Ne

[Ano63-31] Anonymous. News. The American Statistician, 17(5):2–5, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685665.

Anonymous:1963:PNa

[Ano63-32] Anonymous. Personal news. The American Statistician, 17(1):12– 15, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682488.

Anonymous:1963:PNb

[Ano63-33] Anonymous. Personal news. The American Statistician, 17(2): 30–31, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685517.

Anonymous:1963:PNc

[Ano63-34] Anonymous. Personal news. The American Statistician, 17(3): 10–53, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681333.

Anonymous:1963:PNd

[Ano63-35] Anonymous. Personal news. The American Statistician, 17(4): 11–36, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682591.

Anonymous:1963:PNe

[Ano63-36] Anonymous. Personal news. The American Statistician, 17(5):7– 27, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685667.

Anonymous:1963:AMJ

[Ano63-37] Anonymous. The 123rd Annual Meeting: Joint Statistical Meetings. The American Statistician, 17(3):43–49, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681343.

Anonymous:1964:AMJ

[Ano64a] Anonymous. 124th annual meeting: Joint statistical meetings. The American Statistician, 18(4):29–37, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682155.

Anonymous:1964:BMa

[Ano64b] Anonymous. Back matter. The American Statistician, 18(2): 35–36, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682402.

Anonymous:1964:BMb

[Ano64c] Anonymous. Back matter. The American Statistician, 18(3):??, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681486.

Anonymous:1964:BMc

[Ano64d] Anonymous. Back matter. The American Statistician, 18(4): ??, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682159.

Anonymous:1964:BMd

[Ano64e] Anonymous. Back matter. The American Statistician, 18(5):32, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682470.

Anonymous:1964:CNa

[Ano64f] Anonymous. Chapter notes. The American Statistician, 18(2): 32–34, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682401.

Anonymous:1964:CNb

[Ano64g] Anonymous. Chapter notes. The American Statistician, 18(3): 33–36, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681484.

Anonymous:1964:CNc

[Ano64h] Anonymous. Chapter notes. The American Statistician, 18(4): 45–47, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682158.

Anonymous:1964:CNd

[Ano64i] Anonymous. Chapter notes. The American Statistician, 18(5):29– 31, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682469.

Anonymous:1964:ECM

[Ano64j] Anonymous. Edward C. Molina. The American Statistician, 18 (3):36, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681485.

Anonymous:1964:FSAa

[Ano64k] Anonymous. Federal statistical activities. The American Statistician, 18(2):9–10, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682391.

Anonymous:1964:FSAb

[Ano64l] Anonymous. Federal statistical activities. The American Statistician, 18(3):8–9, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681477.

Anonymous:1964:FSAc

[Ano64m] Anonymous. Federal statistical activities. The American Statistician, 18(4):14–48, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682150.

Anonymous:1964:FSAd

[Ano64n] Anonymous. Federal statistical activities. The American Statistician, 18(5):9–10, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682460.

Anonymous:1964:FMa

[Ano64o] Anonymous. Front matter. The American Statistician, 18(1): 1, February 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681604.

Anonymous:1964:FMb

[Ano64p] Anonymous. Front matter. The American Statistician, 18(2): 1–21, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682387.

Anonymous:1964:FMc

[Ano64q] Anonymous. Front matter. The American Statistician, 18(3): 1–35, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681473.

Anonymous:1964:FMd

[Ano64r] Anonymous. Front matter. The American Statistician, 18(4): 1, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682143.

Anonymous:1964:FMe

[Ano64s] Anonymous. Front matter. The American Statistician, 18(5):1– 21, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682456.

Anonymous:1964:GCS

[Ano64t] Anonymous. Government of Canada statistical activities. The American Statistician, 18(4):12–13, October 1964. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682149.

Anonymous:1964:IAS

[Ano64u] Anonymous. Index to the American Statistician August 1947 through December 1963. The American Statistician, 18(1):2– 8, February 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681605.

Anonymous:1964:ISAa

[Ano64v] Anonymous. International statistical activities. The American Statistician, 18(2):8–34, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682390.

Anonymous:1964:ISAb

[Ano64w] Anonymous. International statistical activities. The American Statistician, 18(3):7, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681476.

Anonymous:1964:ISAc

[Ano64x] Anonymous. International statistical activities. The American Statistician, 18(4):12, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682148.

Anonymous:1964:ISAd

[Ano64y] Anonymous. International statistical activities. The American Statistician, 18(5):8–9, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682459.

Anonymous:1964:Na

[Ano64z] Anonymous. News. The American Statistician, 18(2):3–7, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682389.

Anonymous:1964:Nb

[Ano64-27] Anonymous. News. The American Statistician, 18(3):2–6, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681474.

Anonymous:1964:Nc

[Ano64-28] Anonymous. News. The American Statistician, 18(4):3–11, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682145.

Anonymous:1964:Nd

[Ano64-29] Anonymous. News. The American Statistician, 18(5):2–7, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682457.

Anonymous:1964:PNa

[Ano64-30] Anonymous. Personal news. The American Statistician, 18(2): 10–34, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682392.

Anonymous:1964:PNb

[Ano64-31] Anonymous. Personal news. The American Statistician, 18(3): 6–9, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681475.

Anonymous:1964:PNc

[Ano64-32] Anonymous. Personal news. The American Statistician, 18(4): 11–48, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682147.

Anonymous:1964:PNd

[Ano64-33] Anonymous. Personal news. The American Statistician, 18(5):7– 31, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682458.

Anonymous:1964:PRM

[Ano64-34] Anonymous. Program of the Regional Meeting of The Institute of Mathematical Statistics, The Biometric Society (ENAR) with the Section on Physical and Engineering Sciences, ASA. The American Statistician, 18(2):18–19, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682394.

Anonymous:1964:RCP

[Ano64-35] Anonymous. Report of the committee on probability and statistics in the secondary school. The American Statistician, 18(4): 17–25, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682152.

Anonymous:1965:AC

[Ano65a] Anonymous. 125th Anniversary Celebration. The American Statistician, 19(2):18, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682371.

Anonymous:1965:AMJ

[Ano65b] Anonymous. 125th Annual Meeting: Joint Statistical Meetings. The American Statistician, 19(3):26–35, June 1965. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681420.

Anonymous:1965:ATN

[Ano65c] Anonymous. ASA: Then and now: Ten-eighths of a century. The American Statistician, 19(2):29–36, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682375.

Anonymous:1965:BMa

[Ano65d] Anonymous. Back matter. The American Statistician, 19(1): ??, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682333.

Anonymous:1965:BMb

[Ano65e] Anonymous. Back matter. The American Statistician, 19(2): 59–64, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682386.

Anonymous:1965:BMc

[Ano65f] Anonymous. Back matter. The American Statistician, 19(3):48, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681425.

Anonymous:1965:BMd

[Ano65g] Anonymous. Back matter. The American Statistician, 19(4): 48, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681581.

Anonymous:1965:BMe

[Ano65h] Anonymous. Back matter. The American Statistician, 19(5):44, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681902.

Anonymous:1965:CNa

[Ano65i] Anonymous. Chapter notes. The American Statistician, 19(1):38– 40, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682332.

Anonymous:1965:CNb

[Ano65j] Anonymous. Chapter notes. The American Statistician, 19(2): 55–57, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682384.

Anonymous:1965:CNc

[Ano65k] Anonymous. Chapter notes. The American Statistician, 19(3): 45–47, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681424.

Anonymous:1965:CNd

[Ano65l] Anonymous. Chapter notes. The American Statistician, 19(4): 41–43, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681576.

Anonymous:1965:CNe

[Ano65m] Anonymous. Chapter notes. The American Statistician, 19(5):42– 43, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681901.

Anonymous:1965:CG

[Ano65n] Anonymous. Corrado Gini, 1884–1965. The American Statistician, 19(5):40, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681899.

Anonymous:1965:DH

[Ano65o] Anonymous. Dorothy Hinton. The American Statistician, 19(4): 45, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681580.

Anonymous:1965:FSAa

[Ano65p] Anonymous. Federal statistical activities. The American Statistician, 19(1):14–16, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682324.

Anonymous:1965:FSAb

[Ano65q] Anonymous. Federal statistical activities. The American Statistician, 19(2):14–15, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682370.

Anonymous:1965:FSAc

[Ano65r] Anonymous. Federal statistical activities. The American Statistician, 19(3):10–11, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681416.

Anonymous:1965:FSAd

[Ano65s] Anonymous. Federal statistical activities. The American Statistician, 19(4):14–15, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681567.

Anonymous:1965:FSAe

[Ano65t] Anonymous. Federal statistical activities. The American Statistician, 19(5):12–13, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681892.

Anonymous:1965:FMa

[Ano65u] Anonymous. Front matter. The American Statistician, 19(1):1– 37, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682318.

Anonymous:1965:FMb

[Ano65v] Anonymous. Front matter. The American Statistician, 19(2): 1–17, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682366.

Anonymous:1965:FMc

[Ano65w] Anonymous. Front matter. The American Statistician, 19(3): 1–41, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681412.

Anonymous:1965:FMd

[Ano65x] Anonymous. Front matter. The American Statistician, 19(4):1– 38, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681563.

Anonymous:1965:FMe

[Ano65y] Anonymous. Front matter. The American Statistician, 19(5):1– 15, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681888.

Anonymous:1965:GCS

[Ano65z] Anonymous. Government of Canada statistical activities. The American Statistician, 19(4):16–17, October 1965. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681568.

Anonymous:1965:HBB

[Ano65-27] Anonymous. Herman B. Byer. The American Statistician, 19(4): 45, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681579.

Anonymous:1965:ISAa

[Ano65-28] Anonymous. International statistical activities. The American Statistician, 19(1):12–13, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682323.

Anonymous:1965:ISAb

[Ano65-29] Anonymous. International statistical activities. The American Statistician, 19(2):12–15, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682369.

Anonymous:1965:ISAc

[Ano65-30] Anonymous. International statistical activities. The American Statistician, 19(3):9, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681414.

Anonymous:1965:ISAd

[Ano65-31] Anonymous. International statistical activities. The American Statistician, 19(4):12–13, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681565.

Anonymous:1965:ISAe

[Ano65-32] Anonymous. International statistical activities. The American Statistician, 19(5):11–20, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681891.

Anonymous:1965:LKR

[Ano65-33] Anonymous. Lila Knudsen Randolph. The American Statistician, 19(4):45, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681578.

Anonymous:1965:Na

[Ano65-34] Anonymous. News. The American Statistician, 19(1):3–11, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682321.

Anonymous:1965:Nb

[Ano65-35] Anonymous. News. The American Statistician, 19(2):3–15, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682368.

Anonymous:1965:Nc

[Ano65-36] Anonymous. News. The American Statistician, 19(3):2–9, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681413.

Anonymous:1965:Nd

[Ano65-37] Anonymous. News. The American Statistician, 19(4):2–9, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681564.

Anonymous:1965:Ne

[Ano65-38] Anonymous. News. The American Statistician, 19(5):2–10, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681889.

Anonymous:1965:PNa

[Ano65-39] Anonymous. Personal news. The American Statistician, 19(1):11– 35, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682322.

Anonymous:1965:PNb

[Ano65-40] Anonymous. Personal news. The American Statistician, 19(2):54, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682383.

Anonymous:1965:PNc

[Ano65-41] Anonymous. Personal news. The American Statistician, 19(3): 9–14, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681415.

Anonymous:1965:PNd

[Ano65-42] Anonymous. Personal news. The American Statistician, 19(4): 13–47, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681566.

Anonymous:1965:PNe

[Ano65-43] Anonymous. Personal news. The American Statistician, 19(5): 10, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681890.

Anonymous:1965:PIS

[Ano65-44] Anonymous. Program of the International Symposium on Multivariate Analysis. The American Statistician, 19(2):47–48, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682379.

Anonymous:1965:SRM

[Ano65-45] Anonymous. Spring Regional Meetings of The Institute of Mathematical Statistics (Eastern Region), The Biometric Society (ENAR) and The American Statistical Association (Biometrics Section and Section on Physical and Engineering Sciences). The American Statistician, 19(1):30–32, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682330.

Anonymous:1965:WC

[Ano65-46] Anonymous. William S. Connor. The American Statistician, 19 (4):45, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681577.

Anonymous:1966:AMJ

[Ano66a] Anonymous. 126th Annual Meeting: Joint Statistical Meetings. The American Statistician, 20(3):26–36, June 1966. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681497.

Anonymous:1966:AFR

[Ano66b] Anonymous. ASA faculty representatives. The American Statistician, 20(4):6, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685954.

Anonymous:1966:BMa

[Ano66c] Anonymous. Back matter. The American Statistician, 20(1): ??, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681679.

Anonymous:1966:BMb

[Ano66d] Anonymous. Back matter. The American Statistician, 20(2):44, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682720.

Anonymous:1966:BMc

[Ano66e] Anonymous. Back matter. The American Statistician, 20(3):??, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681505.

Anonymous:1966:BMd

[Ano66f] Anonymous. Back matter. The American Statistician, 20(4): 48, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685968.

Anonymous:1966:BMe

[Ano66g] Anonymous. Back matter. The American Statistician, 20(5):??, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682645.

Anonymous:1966:CNa

[Ano66h] Anonymous. Chapter notes. The American Statistician, 20(1):37– 39, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681678.

Anonymous:1966:CNb

[Ano66i] Anonymous. Chapter notes. The American Statistician, 20(2): 39–41, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682717.

Anonymous:1966:CNc

[Ano66j] Anonymous. Chapter notes. The American Statistician, 20(3): 51–54, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681504.

Anonymous:1966:CNd

[Ano66k] Anonymous. Chapter notes. The American Statistician, 20(4): 44–45, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685967.

Anonymous:1966:CNe

[Ano66l] Anonymous. Chapter notes. The American Statistician, 20(5):38– 40, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682644.

Anonymous:1966:CQA

[Ano66m] Anonymous. Correction: Questions and answers. The American Statistician, 20(3):43, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681500.

Anonymous:1966:FRJ

[Ano66n] Anonymous. Faculty representatives — January, 1966. The American Statistician, 20(2):41, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682718.

Anonymous:1966:FID

[Ano66o] Anonymous. Family income distribution statistics published by federal agencies. The American Statistician, 20(1):18–23, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681671.

Anonymous:1966:FSAa

[Ano66p] Anonymous. Federal statistical activities. The American Statistician, 20(1):12–39, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681669.

Anonymous:1966:FSAb

[Ano66q] Anonymous. Federal statistical activities. The American Statistician, 20(2):12–13, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682709.

Anonymous:1966:FSAc

[Ano66r] Anonymous. Federal statistical activities. The American Statistician, 20(3):12–13, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681491.

Anonymous:1966:FSAd

[Ano66s] Anonymous. Federal statistical activities. The American Statistician, 20(4):18–19, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685957.

Anonymous:1966:FMa

[Ano66t] Anonymous. Front matter. The American Statistician, 20(1):1– 6, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681665.

Anonymous:1966:FMb

[Ano66u] Anonymous. Front matter. The American Statistician, 20(2): 1–17, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682705.

Anonymous:1966:FMc

[Ano66v] Anonymous. Front matter. The American Statistician, 20(3): 1–55, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681487.

Anonymous:1966:FMd

[Ano66w] Anonymous. Front matter. The American Statistician, 20(4):1– 40, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685952.

Anonymous:1966:FMe

[Ano66x] Anonymous. Front matter. The American Statistician, 20(5):1– 26, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682632.

Anonymous:1966:GCS

[Ano66y] Anonymous. Government of Canada statistical activities. The American Statistician, 20(5):10–13, December 1966. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682634.

Anonymous:1966:ISAa

[Ano66z] Anonymous. International statistical activities. The American Statistician, 20(1):11–39, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681668.

Anonymous:1966:ISAb

[Ano66-27] Anonymous. International statistical activities. The American Statistician, 20(2):9, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682707.

Anonymous:1966:ISAc

[Ano66-28] Anonymous. International statistical activities. The American Statistician, 20(3):11, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681489.

Anonymous:1966:ISAd

[Ano66-29] Anonymous. International statistical activities. The American Statistician, 20(4):16–17, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685955.

Anonymous:1966:ISAe

[Ano66-30] Anonymous. International statistical activities. The American Statistician, 20(5):14–15, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682635.

Anonymous:1966:Na

[Ano66-31] Anonymous. News. The American Statistician, 20(1):2–10, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681666.

Anonymous:1966:Nb

[Ano66-32] Anonymous. News. The American Statistician, 20(2):2–9, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682706.

Anonymous:1966:Nc

[Ano66-33] Anonymous. News. The American Statistician, 20(3):2–10, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681488.

Anonymous:1966:Nd

[Ano66-34] Anonymous. News. The American Statistician, 20(4):2–15, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685953.

Anonymous:1966:Ne

[Ano66-35] Anonymous. News. The American Statistician, 20(5):2–10, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682633.

Anonymous:1966:PNa

[Ano66-36] Anonymous. Personal news. The American Statistician, 20(1):10– 40, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681667.

Anonymous:1966:PNb

[Ano66-37] Anonymous. Personal news. The American Statistician, 20(2): 9–11, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682708.

Anonymous:1966:PNc

[Ano66-38] Anonymous. Personal news. The American Statistician, 20(3): 11–56, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681490.

Anonymous:1966:PNd

[Ano66-39] Anonymous. Personal news. The American Statistician, 20(4): 17–47, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685956.

Anonymous:1966:PNe

[Ano66-40] Anonymous. Personal news. The American Statistician, 20(5):15– 34, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682636.

Anonymous:1966:PAD

[Ano66-41] Anonymous. Profile-age distribution of ASA membership, 1964. The American Statistician, 20(2):25–28, April 1966. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682713.

Anonymous:1966:PEA

[Ano66-42] Anonymous. Profile-educational attainments of ASA members, 1964. The American Statistician, 20(1):26–28, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681674.

Anonymous:1966:PGC

[Ano66-43] Anonymous. Profile-geographic composition of ASA membership, 1964. The American Statistician, 20(3):37–41, June 1966. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681498.

Anonymous:1966:PSA

[Ano66-44] Anonymous. Profile-statistical activities, methodological techniques, and fields of application of ASA members, 1964. The American Statistician, 20(4):35–37, October 1966. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685964.

Anonymous:1966:SRM

[Ano66-45] Anonymous. Spring Regional Meetings of The Institute of Mathematical Statistics (Eastern Region), The Biometric Society (ENAR) and The American Statistical Association (Biometrics Section and Section on Physical and Engineering Sciences).

The American Statistician, 20(2):34–36, April 1966. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682715.

Anonymous:1967:AMJ

[Ano67a] Anonymous. 127th Annual Meeting: Joint Statistical Meetings. The American Statistician, 21(4):35–44, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682106.

Anonymous:1967:AM

[Ano67b] Anonymous. 1967 Annual Meeting. The American Statistician, 21(1):20, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681909.

Anonymous:1967:BMa

[Ano67c] Anonymous. Back matter. The American Statistician, 21(1): 44, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681918.

Anonymous:1967:BMb

[Ano67d] Anonymous. Back matter. The American Statistician, 21(2): 47–48, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682618.

Anonymous:1967:BMc

[Ano67e] Anonymous. Back matter. The American Statistician, 21(3):??, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682047.

Anonymous:1967:BMd

[Ano67f] Anonymous. Back matter. The American Statistician, 21(4): ??, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682111.

Anonymous:1967:BMe

[Ano67g] Anonymous. Back matter. The American Statistician, 21(5):??, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682665.

Anonymous:1967:CNa

[Ano67h] Anonymous. Chapter notes. The American Statistician, 21(1):42– 43, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681917.

Anonymous:1967:CNb

[Ano67i] Anonymous. Chapter notes. The American Statistician, 21(2): 45–46, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682617.

Anonymous:1967:CNc

[Ano67j] Anonymous. Chapter notes. The American Statistician, 21(3): 45–48, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682046.

Anonymous:1967:CNd

[Ano67k] Anonymous. Chapter notes. The American Statistician, 21(4): 56, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682110.

Anonymous:1967:CNe

[Ano67l] Anonymous. Chapter notes. The American Statistician, 21(5): 36, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682664.

Anonymous:1967:CQA

[Ano67m] Anonymous. Corrigenda: Questions and answers. The American Statistician, 21(4):48, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682108.

Anonymous:1967:FMa

[Ano67n] Anonymous. Front matter. The American Statistician, 21(1):1– 11, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681903.

Anonymous:1967:FMb

[Ano67o] Anonymous. Front matter. The American Statistician, 21(2): 1–11, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682601.

Anonymous:1967:FMc

[Ano67p] Anonymous. Front matter. The American Statistician, 21(3):1–9, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682033.

Anonymous:1967:FMd

[Ano67q] Anonymous. Front matter. The American Statistician, 21(4): 1, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682091.

Anonymous:1967:FMe

[Ano67r] Anonymous. Front matter. The American Statistician, 21(5): ??, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682646.

Anonymous:1967:GP

[Ano67s] Anonymous. Gladys L. Palmer, 1895–1967. The American Statistician, 21(5):35, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682661.

Anonymous:1967:HG

[Ano67t] Anonymous. Herman Guterman. The American Statistician, 21(5):36, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682663.

Anonymous:1967:ISAa

[Ano67u] Anonymous. International statistical activities. The American Statistician, 21(1):8–9, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681906.

Anonymous:1967:ISAb

[Ano67v] Anonymous. International statistical activities. The American Statistician, 21(2):10, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682604.

Anonymous:1967:ISAc

[Ano67w] Anonymous. International statistical activities. The American Statistician, 21(3):8, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682035.

Anonymous:1967:ISAd

[Ano67x] Anonymous. International statistical activities. The American Statistician, 21(4):13–14, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682094.

Anonymous:1967:ISAe

[Ano67y] Anonymous. International statistical activities. The American Statistician, 21(5):9, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682649.

Anonymous:1967:JFD

[Ano67z] Anonymous. James Frederic Dewhurst, 1895–1967. The American Statistician, 21(5):35–36, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682662.

Anonymous:1967:LP

[Ano67-27] Anonymous. Lloyd A. Prochnow. The American Statistician, 21(1):41, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681916.

Anonymous:1967:Na

[Ano67-28] Anonymous. News. The American Statistician, 21(1):2–7, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681904.

Anonymous:1967:Nb

[Ano67-29] Anonymous. News. The American Statistician, 21(2):2–9, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682602.

Anonymous:1967:Nc

[Ano67-30] Anonymous. News. The American Statistician, 21(3):2–7, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682034.

Anonymous:1967:Nd

[Ano67-31] Anonymous. News. The American Statistician, 21(4):4–13, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682093.

Anonymous:1967:Ne

[Ano67-32] Anonymous. News. The American Statistician, 21(5):1–8, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682647.

Anonymous:1967:PNa

[Ano67-33] Anonymous. Personal news. The American Statistician, 21(1):7– 8, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681905.

Anonymous:1967:PNb

[Ano67-34] Anonymous. Personal news. The American Statistician, 21(2): 9–10, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682603.

Anonymous:1967:PNc

[Ano67-35] Anonymous. Personal news. The American Statistician, 21(3): 8–29, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682036.

Anonymous:1967:PNd

[Ano67-36] Anonymous. Personal news. The American Statistician, 21(4): 14–15, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682095.

Anonymous:1967:PNe

[Ano67-37] Anonymous. Personal news. The American Statistician, 21(5): 8, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682648.

Anonymous:1968:AMJ

[Ano68a] Anonymous. 128th Annual Meeting: Joint Statistical Meetings. The American Statistician, 22(3):32–41, June 1968. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681809.

Anonymous:1968:BMa

[Ano68b] Anonymous. Back matter. The American Statistician, 22(1): 40, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681887.

Anonymous:1968:BMb

[Ano68c] Anonymous. Back matter. The American Statistician, 22(2): 39–40, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681993.

Anonymous:1968:BMc

[Ano68d] Anonymous. Back matter. The American Statistician, 22(3):??, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681814.

Anonymous:1968:BMd

[Ano68e] Anonymous. Back matter. The American Statistician, 22(4): ??, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682032.

Anonymous:1968:BMe

[Ano68f] Anonymous. Back matter. The American Statistician, 22(5):??, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681847.

Anonymous:1968:CNa

[Ano68g] Anonymous. Chapter notes. The American Statistician, 22(1):38– 39, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681886.

Anonymous:1968:CNb

[Ano68h] Anonymous. Chapter notes. The American Statistician, 22(2): 36–38, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681992.

Anonymous:1968:CNc

[Ano68i] Anonymous. Chapter notes. The American Statistician, 22(3): 46–48, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681813.

Anonymous:1968:CNd

[Ano68j] Anonymous. Chapter notes. The American Statistician, 22(4): 51–52, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682031.

Anonymous:1968:CNe

[Ano68k] Anonymous. Chapter notes. The American Statistician, 22(5):39– 40, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681846.

Anonymous:1968:CIA

[Ano68l] Anonymous. Correction: An inequality for R/X — and an application to variables sampling. The American Statistician, 22(3):30, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681807. See

[Kir68].

Anonymous:1968:EW

[Ano68m] Anonymous. Edward White, 1875–1968. The American Statistician, 22(3):16, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681796.

Anonymous:1968:ENN

[Ano68n] Anonymous. Erratum notice: “A Note on the Correlation of Parts with Wholes”. The American Statistician, 22(5):21, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681834. See

[BP68].

Anonymous:1968:FMa

[Ano68o] Anonymous. Front matter. The American Statistician, 22(1):1– 17, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681869.

Anonymous:1968:FMb

[Ano68p] Anonymous. Front matter. The American Statistician, 22(2):1, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681975.

Anonymous:1968:FMc

[Ano68q] Anonymous. Front matter. The American Statistician, 22(3):??, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681791.

Anonymous:1968:FMd

[Ano68r] Anonymous. Front matter. The American Statistician, 22(4):1– 49, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682011.

Anonymous:1968:FMe

[Ano68s] Anonymous. Front matter. The American Statistician, 22(5):12– 13, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681829.

Anonymous:1968:ISAa

[Ano68t] Anonymous. International statistical activities. The American Statistician, 22(1):12–13, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681872.

Anonymous:1968:ISAb

[Ano68u] Anonymous. International statistical activities. The American Statistician, 22(2):11–13, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681978.

Anonymous:1968:ISAc

[Ano68v] Anonymous. International statistical activities. The American Statistician, 22(3):15, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681795.

Anonymous:1968:ISAd

[Ano68w] Anonymous. International statistical activities. The American Statistician, 22(4):12, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682014.

Anonymous:1968:Na

[Ano68x] Anonymous. News. The American Statistician, 22(1):2–11, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681870.

Anonymous:1968:Nb

[Ano68y] Anonymous. News. The American Statistician, 22(2):2–10, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681976.

Anonymous:1968:Nc

[Ano68z] Anonymous. News. The American Statistician, 22(3):1–13, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681792.

Anonymous:1968:Nd

[Ano68-27] Anonymous. News. The American Statistician, 22(4):1–11, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682012.

Anonymous:1968:Ne

[Ano68-28] Anonymous. News. The American Statistician, 22(5):1–10, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681830.

Anonymous:1968:PNa

[Ano68-29] Anonymous. Personal news. The American Statistician, 22(1): 11, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681871.

Anonymous:1968:PNb

[Ano68-30] Anonymous. Personal news. The American Statistician, 22(2): 10–15, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681977.

Anonymous:1968:PNc

[Ano68-31] Anonymous. Personal news. The American Statistician, 22(3): 14–16, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681794.

Anonymous:1968:PNd

[Ano68-32] Anonymous. Personal news. The American Statistician, 22(4): 11–45, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682013.

Anonymous:1968:PNe

[Ano68-33] Anonymous. Personal news. The American Statistician, 22(5):10– 11, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681831.

Anonymous:1968:PSI

[Ano68-34] Anonymous. Program of the Second International Symposium on Multivariate Analysis. The American Statistician, 22(2):13–15, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681979.

Anonymous:1969:AMJ

[Ano69a] Anonymous. 129th Annual Meeting: Joint Statistical Meetings. The American Statistician, 23(3):29–42, June 1969. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682579.

Anonymous:1969:BMa

[Ano69b] Anonymous. Back matter. The American Statistician, 23(1):39– 40, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681695.

Anonymous:1969:BMb

[Ano69c] Anonymous. Back matter. The American Statistician, 23(2):??, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681302.

Anonymous:1969:BMc

[Ano69d] Anonymous. Back matter. The American Statistician, 23(3):??, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682585.

Anonymous:1969:BMd

[Ano69e] Anonymous. Back matter. The American Statistician, 23(4): ??, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681749.

Anonymous:1969:BMe

[Ano69f] Anonymous. Back matter. The American Statistician, 23(5):56, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682199.

Anonymous:1969:CNa

[Ano69g] Anonymous. Chapter notes. The American Statistician, 23(1):37– 38, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681694.

Anonymous:1969:CNb

[Ano69h] Anonymous. Chapter notes. The American Statistician, 23(2): 39–40, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681301.

Anonymous:1969:CNc

[Ano69i] Anonymous. Chapter notes. The American Statistician, 23(3): 51–52, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682584.

Anonymous:1969:CNd

[Ano69j] Anonymous. Chapter notes. The American Statistician, 23(4): 56, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681748.

Anonymous:1969:CNe

[Ano69k] Anonymous. Chapter notes. The American Statistician, 23(5):54– 55, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682198.

Anonymous:1969:CPS

[Ano69l] Anonymous. Committee of Presidents of Statistical Societies. The American Statistician, 23(2):37–39, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681300.

Anonymous:1969:CSC

[Ano69m] Anonymous. Conference on statistical computation. The American Statistician, 23(1):31, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681690.

Anonymous:1969:CUS

[Ano69n] Anonymous. Criteria for undertaking surveys for other Federal Government agencies. The American Statistician, 23(1):17–19, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681683.

Anonymous:1969:FAC

[Ano69o] Anonymous. Fiftieth anniversary of census advisory committee. The American Statistician, 23(4):20–22, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681728.

Anonymous:1969:FQS

[Ano69p] Anonymous. First quarter survey of the economic outlook. The American Statistician, 23(2):31–33, April 1969. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681296.

Anonymous:1969:FMa

[Ano69q] Anonymous. Front matter. The American Statistician, 23(1): 1, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681680.

Anonymous:1969:FMb

[Ano69r] Anonymous. Front matter. The American Statistician, 23(2):1, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681285.

Anonymous:1969:FMc

[Ano69s] Anonymous. Front matter. The American Statistician, 23(3): 1–20, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682570.

Anonymous:1969:FMd

[Ano69t] Anonymous. Front matter. The American Statistician, 23(4):1– 50, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681721.

Anonymous:1969:FMe

[Ano69u] Anonymous. Front matter. The American Statistician, 23(5):1– 14, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682175.

Anonymous:1969:ISA

[Ano69v] Anonymous. International statistical activities. The American Statistician, 23(5):12–13, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682178.

Anonymous:1969:Na

[Ano69w] Anonymous. News. The American Statistician, 23(1):2–11, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681681.

Anonymous:1969:Nb

[Ano69x] Anonymous. News. The American Statistician, 23(2):2–8, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681286.

Anonymous:1969:Nc

[Ano69y] Anonymous. News. The American Statistician, 23(3):2–9, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682571.

Anonymous:1969:Nd

[Ano69z] Anonymous. News. The American Statistician, 23(4):1–6, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681722.

Anonymous:1969:Ne

[Ano69-27] Anonymous. News. The American Statistician, 23(5):2–11, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682176.

Anonymous:1969:PNa

[Ano69-28] Anonymous. Personal news. The American Statistician, 23(1): 36, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681693.

Anonymous:1969:PNb

[Ano69-29] Anonymous. Personal news. The American Statistician, 23(2):8, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681287.

Anonymous:1969:PNc

[Ano69-30] Anonymous. Personal news. The American Statistician, 23(3): 9–10, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682572.

Anonymous:1969:PNd

[Ano69-31] Anonymous. Personal news. The American Statistician, 23(4): 6, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681723.

Anonymous:1969:PNe

[Ano69-32] Anonymous. Personal news. The American Statistician, 23(5): 11, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682177.

Anonymous:1969:SQS

[Ano69-33] Anonymous. Second quarter survey of the economic outlook. The American Statistician, 23(3):46–47, June 1969. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682581.

Anonymous:1969:SRM

[Ano69-34] Anonymous. Spring Regional Meetings of The Institute of Mathematical Statistics (Eastern Region), The Biometric Society (ENAR) and The American Statistical Association (Biometrics Section and Section on Physical and Engineering Sciences). The American Statistician, 23(1):29–30, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681689.

Anonymous:1969:SRR

[Ano69-35] Anonymous. Stuart A. Rice/Robert W. Burgess. The American Statistician, 23(3):10, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682573.

Anonymous:1969:TQS

[Ano69-36] Anonymous. Third quarter survey of the economic outlook. The American Statistician, 23(4):45–46, October 1969. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681743.

Anscombe:1961:BS

[Ans61] F. J. Anscombe. Bayesian statistics. The American Statistician, 15(1):21–24, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682504.

Alam:1967:NCC

[AR67] Khursheed Alam and M. Haseeb Rizvi. On non-central chisquared and non-central F distributions. The American Statistician, 21(4):21–22, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682098.

Armstrong:1967:DTM

[Arm67] J. Scott Armstrong. Derivation of theory by means of factor analysis or Tom Swift and his Electric Factor Analysis Machine. The American Statistician, 21(5):17–21, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682655.

Auble:1961:NDB

[Aub61] Arthur G. Auble. A note on the demographic base of household formations in the 60’s. The American Statistician, 15(1):25– 26, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682505.

B:1963:MJM

[B.63] D. M. B. M. Joseph Meehan. The American Statistician, 17(3):54, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681348.

Baba:1960:TSJ

[BA60] Yoshiyuki Baba and Shozo Arita. The teaching of statistics in Japan. The American Statistician, 14(2):13–16, April 1960. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682242.

Bahn:1968:MPT

[Bah68] Anita K. Bahn. Mathematical preparation for training statisticians in biomedical science and public health. The American Statistician, 22(4):18–22, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682017.

Bain:1969:MNN

[Bai69] Lee J. Bain. Moments of a noncentral t and noncentral F - distribution. The American Statistician, 23(4):33–34, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681736.

Bak:1966:ORK

[Bak66] K. Bak. Operations research at the kitchen sink and other applications. The American Statistician, 20(2):15–16, April 1966. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682710.

Bak:1967:ORB

[Bak67] K. Bak. Operations research in the bathtub and other applications. The American Statistician, 21(4):22–24, October 1967. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682099.

Bancroft:1960:OSI

[Ban60a] T. A. Bancroft. The organization of statistical instruction in colleges and universities. The American Statistician, 14(5):14–21, December 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682478.

Bantegui:1960:RSD

[Ban60b] Bernardino G. Bantegui. Recent statistical development in the Philippines. The American Statistician, 14(4):19–32, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681379.

Bancroft:1961:POP

[Ban61] T. A. Bancroft. Practical organizational plans for a complete university computing service. The American Statistician, 15(5):10– 18, December 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681612.

Barberi:1960:TSI

[Bar60] Benedetto Barberi. On the teaching of statistics in Italy. The American Statistician, 14(2):17–18, April 1960. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682243.

Brown:1964:LE

[BCU^{+}64] Robert H. Brown, Joseph T. Crymes, Hans Ury, G. G. Eldredge, L. J. Jacobson, and Derrick G. Low Choy. Letters to the Editor. The American Statistician, 18(5):28–31, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682468.

Beach:1960:TSC

[Bea60] E. F. Beach. The teaching of statistics in Canada. The American Statistician, 14(5):17–18, December 1960. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682479.

Bean:1967:CWA

[Bea67a] Louis H. Bean. Crops, weather, and the agricultural revolution. The American Statistician, 21(3):10–14, June 1967. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682037.

Bean:1967:RMB

[Bea67b] Louis H. Bean. Rejoinder to “Mr. Bean and the Weather”. The American Statistician, 21(5):13–14, December 1967. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682652. See [Sha67].

Behboodian:1967:EAV

[Beh67] Javad Behboodian. On the expectation of the absolute value of a random vector. The American Statistician, 21(4):25–26, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682101.

Berkson:1963:SLC

[Ber63] Joseph Berkson. Smoking and lung cancer. The American Statistician, 17(4):15–22, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682593.

Barnett:1968:OSA

[BGB68] E. H. Barnett, G. E. Gautney, and Ernest E. Blanche. Opportunities for statisticians in the area of computers. The American Statistician, 22(4):46–47, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682029.

Bancroft:1968:TRM

[BH68] T. A. Bancroft and Roy D. Hickman. Thesis requirements for the Masters Degree in Statistics at American universities. The American Statistician, 22(4):28–30, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682019.

Bross:1965:LE

[BHD^{+}65] Irwin D. J. Bross, J. A. Hartigan, John M. Dwyer, D. M. Olsson, Thomas M. Karunas, Max Halperin, and John E. Freund. Letters to the Editor. The American Statistician, 19(4):39–40, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681575.

Brunk:1968:CBA

[BHW68] H. D. Brunk, James E. Holstein, and Frederick Williams. A comparison of binomial approximations to the hypergeometric distribution. The American Statistician, 22(1):24–26, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681878.

Birnbaum:1962:AVF

[Bir62] Allan Birnbaum. Another view on the foundations of statistics. The American Statistician, 16(1):17–21, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681948.

Bowman:1961:FSA

[BKH61] Raymond T. Bowman, James W. Knowles, and Ray Hurley. Federal statistical activities. The American Statistician, 15(1):9– 15, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682502.

Blanche:1968:OSA

[Bla68] Ernest E. Blanche. Opportunities for statisticians in the area of computers. The American Statistician, 22(3):44–45, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681811.

Bliss:1969:CBB

[Bli69] C. I. Bliss. Communication between biologists and statisticians, A case study. The American Statistician, 23(4):15–20, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681727.

Bowman:1962:SRU

[BM62] Raymond T. Bowman and Margaret E. Martin. Special report on unemployment statistics: Meaning and measurement. The American Statistician, 16(4):14–22, October 1962. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681432.

Bowman:1964:FSA

[BM64] R. T. Bowman and Oskar Morgenstern. Fide sed ante vide. The American Statistician, 18(4):15–25, October 1964. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682151.

Bolch:1968:MUE

[Bol68] Ben W. Bolch. More on unbiased estimation of the standard deviation. The American Statistician, 22(3):27, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681802. See

[Mar68a, Cur68b, Bol68, Cur68a].

Bondy:1969:TEH

[Bon69] Warren H. Bondy. A test of an experimental hypothesis of negligible difference between means. The American Statistician, 23 (5):28–30, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682186.

Bowman:1963:PCa

[Bow63a] Raymond T. Bowman. The President’s column. The American Statistician, 17(1):1–32, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682485.

Bowman:1963:PCb

[Bow63b] Raymond T. Bowman. The President’s column. The American Statistician, 17(2):1–35, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685508.

Bowman:1963:PCc

[Bow63c] Raymond T. Bowman. The President’s column. The American Statistician, 17(3):1–26, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681331.

Bowman:1963:PCd

[Bow63d] Raymond T. Bowman. The President’s column. The American Statistician, 17(4):1–24, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682587.

Bowman:1963:PCe

[Bow63e] Raymond T. Bowman. The President’s column. The American Statistician, 17(5):1–15, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685664.

Bowker:1964:PCa

[Bow64a] Albert H. Bowker. The President’s column. The American Statistician, 18(2):2, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682388.

Bowker:1964:PCb

[Bow64b] Albert H. Bowker. The President’s column. The American Statistician, 18(4):2, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682144.

Bowman:1964:CQN

[Bow64c] Raymond T. Bowman. Comments on “Qui Numerare Incipit Errare Incipit” by Oskar Morgenstern. The American Statistician, 18 (3):10–20, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681478.

Bartko:1968:NCP

[BP68] John J. Bartko and Karen D. Pettigrew. A note on the correlation of parts with wholes. The American Statistician, 22(4):41, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682026. See

erratum [Ano68n].

Bradley:1963:ROC

[Bra63] James V. Bradley. Rank order correlation scatter diagrams without plotting points. The American Statistician, 17(2):14–15, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685513.

Bradley:1966:FW

[Bra66] Ralph A. Bradley. Frank Wilcoxon. The American Statistician, 20(1):32–33, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681676.

Braverman:1967:SSC

[Bra67] Shirley Braverman. Sample size comparisons based on functions of order statistics. The American Statistician, 21(5):22–24, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682657.

Bross:1968:LE

[BRG^{+}68] Irwin D. J. Bross, Donald Richter, John Gurland, W. Duane Evans, and T. A. Bancroft. Letters to the Editor. The American Statistician, 22(5):37–38, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681845.

Bross:1961:SDC

[Bro61] Irwin D. J. Bross. Statistical dogma: A challenge. The American Statistician, 15(3):14–15, June 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682625.

Bross:1963:LAS

[Bro63] Irwin D. J. Bross. Linguistic analysis of a statistical controversy. The American Statistician, 17(1):18–21, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682490.

Bross:1964:CIR

[Bro64a] Irwin D. J. Bross. Correction: Is it really wrong? The American Statistician, 18(4):11, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682146.

Bross:1964:IRW

[Bro64b] Irwin D. J. Bross. Is it really wrong? The American Statistician, 18(3):21–22, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681480.

Brugger:1969:NUE

[Bru69] Richard M. Brugger. A note on unbiased estimation of the standard deviation. The American Statistician, 23(4):32, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681734.

Burdick:1960:SI

[BS60] E. Douglass Burdick and K. C. Seal. Statistics in india. The American Statistician, 14(1):23–25, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682130.

Bowman:1960:FSA

[BTGK60] Raymond T. Bowman, Conrad Taeuber, Clayton Gehman, and Frank S. Kristof. Federal statistical activities. The American Statistician, 14(1):10–14, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682126.

Burns:1962:SRE

[Bur62] Arthur F. Burns. Some reflections on the employment act. The American Statistician, 16(5):10–18, December 1962. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681959.

Canner:1969:SCR

[Can69] Paul L. Canner. Some curious results using minimum variance linear unbiased estimators. The American Statistician, 23(5):39– 40, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682191.

Carmer:1969:CDG

[CC69] S. G. Carmer and F. B. Cady. Computerized data generation for teaching statistics. The American Statistician, 23(5):33–35, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682188.

Clelland:1963:DC

[CFG63] Richard Clelland, Denis A. Flagg, and Gerald J. Glasser. Discussants’ comments. The American Statistician, 17(3):24–26, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681338.

Chambers:1968:BWP

[Cha68] John M. Chambers. A British Working Party on Statistical Computing. The American Statistician, 22(2):19–20, April 1968. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681981.

Chew:1968:SUA

[Che68] Victor Chew. Some useful alternatives to the normal distribution. The American Statistician, 22(3):22–24, June 1968. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681799.

Chiang:1966:ERR

[Chi66] C. L. Chiang. On the expectation of the reciprocal of a random variable. The American Statistician, 20(4):28, October 1966. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685960.

Costis:1963:SEG

[CL63] Harry G. Costis and P. John Lymberopoulos. Statistical education in Greece. The American Statistician, 17(4):27, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682595.

Clague:1968:JP

[Cla68] Ewan Clague. Jacob Perlman, 1899–1968. The American Statistician, 22(3):45, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681812.

Cohen:1968:SIS

[Coh68] Wilbur J. Cohen. Social indicators: Statistics for public policy. The American Statistician, 22(4):14–16, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682015.

Cohen:1969:DSF

[Coh69] Samuel E. Cohen. A device for sampling from finite populations on successive occasions. The American Statistician, 23(4):26–27, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681730.

Collins:1965:ARM

[Col65] John F. Collins. Abstract of remarks by Mayor John F. Collins of Boston. The American Statistician, 19(2):18–19, April 1965. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682372.

Conway:1965:RNM

[Con65] Lawrence V. Conway. Report of National Membership Committee. The American Statistician, 19(3):40, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681422.

Corcoran:1963:CSC

[Cor63a] Thomas F. Corcoran. On the confidential status of census reports. The American Statistician, 17(3):33–40, June 1963. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681341.

Cornfield:1963:HFD

[Cor63b] Jerome Cornfield. Harold Fred Dorn. The American Statistician, 17(3):53, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681345.

Cornfield:1966:STS

[Cor66] Jerome Cornfield. Sequential trials, sequential analysis and the likelihood principle. The American Statistician, 20(2):18–23, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682711.

Cox:1962:HMS

[Cox62] Edwin B. Cox. Henry Moore and the “Statistical complement of pure economics”. The American Statistician, 16(2):10–13, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681536.

Cox:1965:TSB

[Cox65] Edwin B. Cox. The teaching of statistics in business schools — a summary report. The American Statistician, 19(1):17–19, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682326.

Cox:1968:SAE

[Cox68] C. Philip Cox. A statistical application of Euler’s Theorem. The American Statistician, 22(2):29, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681987.

Craig:1961:FCM

[Cra61] C. C. Craig. The first course in mathematical statistics. The American Statistician, 15(5):14–25, December 1961. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681613.

Cureton:1964:X

[Cur64] Edward E. Cureton. Z, t, F or β ? The American Statistician, 18(2):26–31, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682398.

Cureton:1968:PCU

[Cur68a] Edward E. Cureton. Priority correction to “Unbiased Estimation of the Standard Deviation”. The American Statistician, 22(3):27, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681803. See

[Mar68a, Cur68b, Bol68].

Cureton:1968:UES

[Cur68b] Edward E. Cureton. Unbiased estimation of the standard deviation. The American Statistician, 22(1):22, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681876. See

[Bol68, Cur68a].

Darlington:1969:DLS

[Dar69] Richard B. Darlington. Deriving least-squares weights without calculus. The American Statistician, 23(5):41–42, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682193.

Deming:1962:LE

[Dem62] W. Edward Deming. Letter to the Editor. The American Statistician, 16(1):16, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681947.

Deming:1965:KK

[Dem65] W. Edwards Deming. Kenichi Koyanagi. The American Statistician, 19(2):53, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682382.

Deming:1967:WS

[Dem67] W. Edwards Deming. Walter A. Shewhart, 1891–1967. The American Statistician, 21(2):39–40, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682615.

Davis:1965:RSC

[DFSM65] Joseph S. Davis, Lester R. Frankel, Frederick F. Stephan, and Geoffrey H. Moore. Report of the subcommittee on criteria and principles. The American Statistician, 19(1):22–27, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682328.

Dyar:1969:CCH

[DG69] Robert Dyar and William R. Gaffey. Current challenges to health statisticians. The American Statistician, 23(5):19–22, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682181.

Densen:1964:OAP

[DS64] Paul M. Densen and Hyman Sardy. Official agencies and problems of recruitment to the statistical field. The American Statistician, 18(3):23–26, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681481.

Dubey:1965:SSC

[Dub65] Satya D. Dubey. Statistical solutions of a combinatorial problem. The American Statistician, 19(4):30–31, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681572.

Dubey:1966:GTD

[Dub66] Satya D. Dubey. Graphical tests for discrete distributions. The American Statistician, 20(3):23–24, June 1966. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681494.

Duchan:1969:RBS

[Duc69] Alan I. Duchan. A relationship between the F and t statistics and the simple correlation coefficients in classical least squares regression. The American Statistician, 23(3):27–28, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682578.

Dunn:1967:IND

[Dun67] Edgar S. Dunn, Jr. The idea of a National Data Center and the issue of personal privacy. The American Statistician, 21(1):21– 27, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681910.

Dutton:1963:LC

[Dut63] Arthur M. Dutton. S. Lee Crump. The American Statistician, 17 (3):54, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681347.

Eisenhart:1967:AISb

[EB67] Churchill Eisenhart and Allan Birnbaum. Anniversaries in 1966– 67 of interest to statisticians. The American Statistician, 21(3): 22–29, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682040.

Eckler:1969:RWB

[Eck69] A. Ross Eckler. Robert Wilbur Burgess, 1887–1969. The American Statistician, 23(4):48–49, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681745.

Edgington:1963:JAD

[Edg63] Eugene S. Edgington. Joint analysis of differences in central tendency and variability. The American Statistician, 17(3):28–30, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681339.

Edwards:1969:RBT

[Edw69] John B. Edwards. The relation between the F -test and R̄^{2}. The American Statistician, 23(5):28, December 1969. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682185.

Eldridge:1963:LE

[EH63] Paul Eldridge and H. Leon Harter. Letter to the Editor. The American Statistician, 17(1):21–32, February 1963. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682491.

Eisenhart:1965:AIS

[Eis65] Churchill Eisenhart. Anniversaries in 1965 of interest to statisticians. The American Statistician, 19(5):21–29, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681894.

Eisenhart:1967:AISa

[Eis67] Churchill Eisenhart. Anniversaries in 1966–67 of interest to statisticians. The American Statistician, 21(2):32–34, April 1967. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682613.

Elston:1969:AFT

[Els69] R. C. Elston. An analogue to Fieller’s theorem using Scheffé’s solution to the Fisher–Behrens problem. The American Statistician, 23(1):26–28, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681688.

Ferber:1966:ASC

[Fer66] Robert Ferber. Anticipations statistics and consumer behavior. The American Statistician, 20(4):20–24, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685958.

Ferber:1967:ASC

[Fer67] Robert Ferber. Anticipations statistics and consumer behavior: A rejoinder. The American Statistician, 21(2):14, April 1967. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682606.

Fleiss:1966:NER

[Fle66] Joseph L. Fleiss. A note on the expectation of the reciprocal of a random variable. The American Statistician, 20(1):25, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681673.

F:1967:LE

[FNG^{+}67] J. E. F., Gottfried E. Noether, N. T. Gridgeman, Lawrence Harris, A. F. Goodman, Ernest M. Scheuer, and H. Fairfield Smith. Letters to the Editor. The American Statistician, 21(2):41–44, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682616.

Freeman:1963:TSD

[Fre63a] William W. K. Freeman. Training of statisticians in diplomacy to maintain their integrity. The American Statistician, 17(5):16– 20, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685671.

Freund:1963:WRE

[Fre63b] R. J. Freund. A warning of roundoff errors in regression. The American Statistician, 17(5):13–15, December 1963. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685670.

Freund:1965:PP

[Fre65] John E. Freund. Puzzle or paradox? The American Statistician, 19(4):29–44, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681571.

Freundenthal:1967:EJG

[Fre67] Alfred M. Freundenthal. Emil J. Gumbel. The American Statistician, 21(1):41, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681915.

Fattu:1964:BV

[FU64] N. A. Fattu and V. R. Rao Uppuluri. On a bias in voting. The American Statistician, 18(5):19–20, December 1964. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682464.

Freeman:1964:SWP

[FW64] Linton C. Freeman and Roland Werner. Statistical word play. The American Statistician, 18(5):18, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682463.

Gainsbrugh:1961:PCa

[Gai61a] Martin R. Gainsbrugh. The President’s column. The American Statistician, 15(2):1–28, April 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682722.

Gainsbrugh:1961:PCb

[Gai61b] Martin R. Gainsbrugh. The President’s column. The American Statistician, 15(3):1–20, June 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682620.

Gainsbrugh:1961:PCc

[Gai61c] Martin R. Gainsbrugh. The President’s column. The American Statistician, 15(4):1–42, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681816.

Gainsbrugh:1961:PCd

[Gai61d] Martin R. Gainsbrugh. The President’s column. The American Statistician, 15(5):1–21, December 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681607.

Gardner:1968:NEV

[Gar68] John W. Gardner. No easy victories. The American Statistician, 22(1):14–16, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681873.

Good:1969:LE

[GBC69] I. J. Good, Richard M. Brugger, and W. P. Cooke. Letters to the Editor. The American Statistician, 23(1):35, February 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681692.

Grier:1968:EKM

[GD68] J. Brown Grier and Raymond Ditrichs. The estimation of knowledge by multiple-choice tests. The American Statistician, 22(4): 35–36, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682022.

Geary:1966:NRH

[Gea66] R. C. Geary. A note on residual heterovariance and estimation efficiency in regression. The American Statistician, 20(4):30–31, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685962.

Gehan:1968:NBP

[Geh68] Edmund A. Gehan. Note on the “Birthday Problem”. The American Statistician, 22(2):28, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681986. See [Nau68].

Grove:1966:LE

[GFB^{+}66] Ernest Wilson Grove, John E. Freund, D. S. Burdick, Stanley L. Payne, J. C. Koop, G. Calot, Edward E. Cureton, C. Philip Cox, and Sherwood C. Frey, Jr. Letters to the Editor. The American Statistician, 20(3):47–49, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681502.

Goldsmith:1965:LE

[GFG^{+}65] Charles H. Goldsmith, Joseph L. Fleiss, A. F. Goodman, Gene Laska, Henry Tucker, I. J. Good, Paul Eldridge, Michael F. Capobianco, N. F. Laubscher, and Ira Marshak. Letters to the Editor. The American Statistician, 19(3):42–44, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681423.

Gibbons:1968:MEE

[Gib68] Jean D. Gibbons. Mutually exclusive events, independence and zero correlation. The American Statistician, 22(5):31–32, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681841.

Gittelsohn:1969:OP

[Git69] Alan M. Gittelsohn. An occupancy problem. The American Statistician, 23(2):11–12, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681289.

Graham:1968:LE

[GKH^{+}68] Jack E. Graham, Kern O. Kymn, R. J. Hader, Jaakko K. Kihlberg, Jay H. Herson, and William E. Schotz. Letters to the Editor. The American Statistician, 22(2):35, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681991.

Gross:1969:LE

[GKK^{+}69] Alan J. Gross, William F. Kwolek, R. N. Kulkarni, Daniel Zelinsky, John P. Hoyt, James K. Brewer, William Knight, Nathan Mantel, C. Mitchell Dayton, W. P. Cooke, Harry D. Levine, and Robert F. Thimmesh. Letters to the Editor. The American Statistician, 23(5):49–53, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682197.

Geary:1968:STM

[GL68] R. C. Geary and C. E. V. Leser. Significance tests in multiple regression. The American Statistician, 22(1):20–21, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681875.

Greenhouse:1966:MBS

[GMN66] Samuel W. Greenhouse, Margaret P. Martin, and Horace W. Norton. To members of the Biometrics Section, ASA. The American Statistician, 20(3):44–46, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681501.

Goldman:1961:LE

[Gol61] Thomas A. Goldman. Letter to the Editor. The American Statistician, 15(1):36, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682510.

Good:1965:LE

[Goo65] I. J. Good. Letter to the Editor. The American Statistician, 19 (3):43, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681423.

Goodman:1968:ICS

[Goo68] A. F. Goodman. The interface of computer science and statistics: An historical perspective. The American Statistician, 22(3):17–20, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681797.

Good:1969:HRR

[Goo69a] I. J. Good. How random are random numbers? The American Statistician, 23(4):42–45, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681742.

Goodman:1969:ISD

[Goo69b] Michael L. Goodman. On the incidence of swept double-headers. The American Statistician, 23(5):15–17, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682179.

Gordon:1964:PPV

[Gor64] Charles K. Gordon, Jr. On the probability of past vs. future events. The American Statistician, 18(2):20–24, April 1964. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682395.

Grizzle:1967:CCT

[Gri67] James E. Grizzle. Continuity correction in the χ^{2}-test for 2×2 tables. The American Statistician, 21(4):28–32, October 1967. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682103.

Gridgeman:1968:PS

[Gri68] N. T. Gridgeman. Probability and sex. The American Statistician, 22(3):29, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681805.

Glaser:1967:DFS

[GRW67] E. Glaser, D. Rosenblatt, and M. K. Wood. The design of a Federal Statistical Data Center. The American Statistician, 21 (1):12–20, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681908.

Gani:1965:BJA

[GS65] J. Gani and A. Spier. The birth of the journal of applied probability. The American Statistician, 19(4):18–36, October 1965. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681569.

Gurney:1969:MB

[GT69] Margaret Gurney and Benjamin F. Tepping. Max A. Bershad, 1913–1969. The American Statistician, 23(4):49, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681746.

Guenther:1969:SCI

[Gue69] William C. Guenther. Shortest confidence intervals. The American Statistician, 23(1):22–25, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681686.

Gurland:1967:ISE

[Gur67] John Gurland. An inequality satisfied by the expectation of the reciprocal of a random variable. The American Statistician, 21(2): 24–25, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682610.

Gurland:1968:IER

[Gur68] John Gurland. Inequalities of expectations of random variables derived by monotonicity or convexity. The American Statistician, 22(2):26–27, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681985.

Hamdan:1969:NBP

[HAB69] M. A. Hamdan and H. A. Al-Bayyati. A note on the bivariate Poisson distribution. The American Statistician, 23(4):32–33, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681735.

Hader:1967:RRP

[Had67] R. J. Hader. Random roommate pairing of Negro and White students. The American Statistician, 21(5):24–26, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682658.

Hahn:1968:SUT

[Hah68] Gerald J. Hahn. Some uses of time-sharing computers in statistical analysis. The American Statistician, 22(5):14–18, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681832.

Haitovsky:1966:NRP

[Hai66] Yoel Haitovsky. A note on regression on principal components. The American Statistician, 20(4):28–29, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685961.

Haitovsky:1969:NMR

[Hai69] Yoel Haitovsky. A note on the maximization of R^{2}. The American Statistician, 23(1):20–21, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681684.

Hamburg:1961:EN

[Ham61] Morris Hamburg. Editor’s note. The American Statistician, 15 (1):1, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682498.

Hamburg:1963:JPB

[Ham63] Morris Hamburg. J. Parker Bursk. The American Statistician, 17 (3):53–54, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681346.

Hansen:1960:PCa

[Han60a] Morris H. Hansen. The President’s column. The American Statistician, 14(1):1–38, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682123.

Hansen:1960:PCb

[Han60b] Morris H. Hansen. The President’s column. The American Statistician, 14(2):1–34, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682238.

Hansen:1960:PCc

[Han60c] Morris H. Hansen. The President’s column. The American Statistician, 14(3):1–45, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682161.

Hansen:1960:PCd

[Han60d] Morris H. Hansen. The President’s column. The American Statistician, 14(4):1–41, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681374.

Hansen:1960:PCe

[Han60e] Morris H. Hansen. The President’s column. The American Statistician, 14(5):1–26, December 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682472.

Hanurav:1968:SRA

[Han68] T. V. Hanurav. Some remarks on Ajgaonkar’s paper “The Effect of Increasing Sample Size on the Precision of an Estimator”. The American Statistician, 22(4):39–40, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682025. See [Ajg67].

Harter:1962:SAF

[Har62] H. Leon Harter. Statistics in the Air Force Research Program. The American Statistician, 16(4):23–24, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681433.

Hartley:1963:DBC

[Har63] H. O. Hartley. In Dr. Bayes’ consulting Room. The American Statistician, 17(1):22–24, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682492.

Harris:1965:NDB

[Har65] Lawrence Harris. A note on the distinction between enumerative and analytical surveys for business decisions. The American Statistician, 19(5):29, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681895.

Hauser:1960:SAS

[Hau60] Philip M. Hauser. Samuel Andrew Stouffer, 1900–1960. The American Statistician, 14(4):36, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681384.

Hauser:1962:PCa

[Hau62a] Philip M. Hauser. The President’s column. The American Statistician, 16(1):1–32, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681941.

Hauser:1962:PCb

[Hau62b] Philip M. Hauser. The President’s column. The American Statistician, 16(2):1–24, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681531.

Hauser:1962:PCc

[Hau62c] Philip M. Hauser. The President’s column. The American Statistician, 16(3):1–46, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682064.

Hauser:1962:PCd

[Hau62d] Philip M. Hauser. The President’s column. The American Statistician, 16(4):1–22, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681427.

Hauser:1962:PCe

[Hau62e] Philip M. Hauser. The President’s column. The American Statistician, 16(5):1–35, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681953.

Hauser:1968:MT

[Hau68] Philip M. Hauser. Martin Taitel. The American Statistician, 22 (1):30–36, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681881.

Houseman:1967:CPM

[HB67] Earl E. Houseman and Joseph A. Becker. A centenary profile of methods for agricultural surveys. The American Statistician, 21 (2):15–21, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682607.

Hagans:1963:LE

[HDC63] James A. Hagans, Lawrence Danziger, and Jerome Cornfield. Letters to the Editor. The American Statistician, 17(5):21–24, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685672.

Herbach:1963:MN

[Her63] Leon H. Herbach. Monroe L. Norden. The American Statistician, 17(3):54, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681349.

Harter:1961:LE

[HGH61] H. Leon Harter, Paul C. Glick, and David M. Heer. Letter to the Editor. The American Statistician, 15(3):21–27, June 1961. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682627.

Halperin:1965:RSS

[HHH65] Max Halperin, H. O. Hartley, and P. G. Hoel. Recommended standards for statistical symbols and notation. COPSS Committee on Symbols and Notation. The American Statistician, 19(3):12–14, June 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681417.

Hoffmann:1961:TS

[Hof61] Robert G. Hoffmann. On teaching statistics. The American Statistician, 15(5):17–18, December 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681614.

Hogben:1968:DSV

[Hog68] David Hogben. The distribution of the sample variance from a two-point binomial population. The American Statistician, 22(5): 30, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681840.

Hogg:1969:SOU

[Hog69] Robert V. Hogg. Some observations on undergraduate programs for statistics majors. The American Statistician, 23(3):21–24, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682575.

Horowitz:1960:CTA

[Hor60] C. Morris Horowitz. Card tabulation by adding machine. The American Statistician, 14(2):21–25, April 1960. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682245.

Hoyt:1968:SAS

[Hoy68] John P. Hoyt. A simple approximation to the standard normal probability density function. The American Statistician, 22(2): 25–26, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681984.

Hoyt:1969:TIE

[Hoy69] John P. Hoyt. Two instructive examples of maximum likelihood estimates. The American Statistician, 23(2):14, April 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681291.

Hartley:1963:LE

[HPWQ63] H. O. Hartley, M. C. Pike, Edwin B. Wilson, and Marian W. Quirk. Letters to the Editor. The American Statistician, 17(4): 31–32, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682597.

Hunter:1967:LE

[HSG^{+}67] J. Stuart Hunter, David Salsburg, Ernest Wilson Grove, Irwin D. J. Bross, and Gabriel Alon. Letters to the Editor. The American Statistician, 21(1):9–40, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681907.

Hudson:1968:SCM

[Hud68] Derek J. Hudson. A short cut method for estimating only one of two parameters from a set of order statistics. The American Statistician, 22(5):23–24, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681836.

Hultquist:1966:DTR

[Hul66] Robert Hultquist. Diagramming theorems relative to quadratic forms. The American Statistician, 20(5):31, December 1966. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682641.

Ifram:1969:NST

[Ifr69] Adnan F. Ifram. A note on sequential tests. The American Statistician, 23(4):36–37, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681738.

Ives:1967:CMN

[IG67] Kenneth H. Ives and Jean D. Gibbons. A correlation measure for nominal data. The American Statistician, 21(5):16–17, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682654.

Ipsen:1965:SHM

[Ips65] Johannes Ipsen. Statistical hurdles in the medical career. The American Statistician, 19(3):22–24, June 1965. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681419.

Jaech:1966:AAM

[Jae66] John L. Jaech. An alternate approach to missing value estimation. The American Statistician, 20(5):27–29, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682639.

Jaffe:1960:NCM

[Jaf60] A. J. Jaffe. A note on “Characteristics of the Membership of the American Statistical Association, 1958”. The American Statistician, 14(1):27–31, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682131.

Jarrett:1968:MEH

[Jar68] Rheem F. Jarrett. A minor exercise in history. The American Statistician, 22(3):25–26, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681800.

Jen:1967:PT

[JF67] Frank C. Jen and Walter Feibes. A permutation theorem. The American Statistician, 21(3):31, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682043.

Johnson:1967:MLD

[JH67] Arthur F. Johnson and Edward Hines, Jr. Minimum location design for the r th order polynomial in k variables. The American Statistician, 21(3):30, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682041.

Johnson:1965:NCI

[Joh65] M. Clemens Johnson. Note on the computer as an instructional tool in statistics. The American Statistician, 19(4):32–36, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681573.

Johnson:1968:FM

[Joh68] Robert E. Johnson. Francis E. McIntyre. The American Statistician, 22(1):36–37, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681885.

Joiner:1968:MDR

[JR68] Brian L. Joiner and Joan R. Rosenblatt. The mean deviation and range for n = 3. The American Statistician, 22(4):37–38, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682024.

Karon:1964:NTA

[Kar64] Bertram P. Karon. A note on the treatment of age as a variable in regression equations. The American Statistician, 18(3):27–28, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681482.

Katona:1967:ASC

[Kat67] George Katona. Anticipations statistics and consumer behavior. The American Statistician, 21(2):12–13, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682605.

Knowles:1961:FSA

[KCSP61] James W. Knowles, Tom Corcoran, Jane Samuels, and Jacob Perlman. Federal statistical activities. The American Statistician, 15 (2):11–13, April 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682726.

King:1961:LE

[KDMS61] E. P. King, W. Edwards Deming, Leo Mandel, and W. H. Somermeijer. Letters to the Editor. The American Statistician, 15(4): 12–23, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681822.

Kenadjian:1965:TDE

[Ken65] Berdj Kenadjian. A tribute to Dr. Ernest J. Engquist. The American Statistician, 19(1):2–27, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682320.

Keppel:1963:ESE

[Kep63] Francis Keppel. Educational statistics and educational policy. The American Statistician, 17(5):9–12, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685669.

Kerrich:1961:RR

[Ker61] J. E. Kerrich. Random remarks. The American Statistician, 15 (3):16–20, June 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682626.

Kerrich:1966:FLW

[Ker66] J. E. Kerrich. Fitting the line y = ax when errors of observation are present in both variables. The American Statistician, 20(1):24, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681672.

Kerrich:1964:LE

[KGvdB^{+}64] J. E. Kerrich, N. T. Gridgeman, Anthony R. van der Burg, Bert van Uitert, Robert W. Hiorns, Gilbert M. Oster, Walter D. Fisher, Ed Capen, and Hugh J. Miser. Letters to the Editor. The American Statistician, 18(4):41–44, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682157.

Kaysen:1969:RTF

[KHH^{+}69] Carl Kaysen, Charles C. Holt, Richard Holton, George Kozmetsky, H. Russell Morrison, and Richard Ruggles. Report of the task force on the storage of and access to government statistics. The American Statistician, 23(3):11–19, June 1969. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682574.

King:1966:CFS

[Kin66] Arnold J. King. Charles F. Sarle. The American Statistician, 20 (4):43, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685966.

Kirkpatrick:1968:IAV

[Kir68] R. L. Kirkpatrick. An inequality for R/X — and an application to variables sampling. The American Statistician, 22(2):29–30, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681988. See

correction [Ano68l].

Kish:1962:LE

[KJ62] Leslie Kish and Denis F. Johnston. Letters to the Editor. The American Statistician, 16(3):39, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682074.

Klotz:1969:SPS

[Klo69] Jerome Klotz. A simple proof of Scheffé’s multiple comparison theorem for contrasts in the one-way layout. The American Statistician, 23(5):44–45, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682195.

Knapp:1962:CUR

[Kna62] Thomas R. Knapp. Classroom use of rank-order data. The American Statistician, 16(4):25–26, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681434.

Koop:1968:ERE

[Koo68] J. C. Koop. An exercise in ratio estimation. The American Statistician, 22(1):29–30, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681880.

Kotz:1965:STR

[Kot65] Samuel Kotz. Statistical terminology — Russian vs. English — in the light of the development of statistics in the U.S.S.R. The American Statistician, 19(3):16–22, June 1965. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681418.

Kramer:1966:SRM

[Kra66] Clyde Young Kramer. Simple regression in multiple regression notation. The American Statistician, 20(3):25, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681496.

Kruskal:1968:NNC

[Kru68] William Kruskal. Note on a note by C. S. Pillai. The American Statistician, 22(5):24–25, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681837. See [Pil68].

Kullback:1968:ATS

[Kul68] S. Kullback. An application of tables of symmetric functions to certain distributions. The American Statistician, 22(5):25–27, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681838.

Kurtz:1968:RTS

[Kur68] Thomas E. Kurtz. A role of time-sharing computing in statistical research. The American Statistician, 22(5):19–21, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681833.

Kullback:1967:PSI

[KY67] S. Kullback and H. Y. Yasaimaibodi. A property of symmetric (interchangeable) random variables. The American Statistician, 21 (1):32–33, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681913.

Kymn:1968:LE

[Kym68] Kern O. Kymn. Letters to the Editor. The American Statistician, 22(1):36, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681884.

Laubscher:1965:IT

[Lau65] Nico F. Laubscher. Interpolation in F -tables. The American Statistician, 19(1):28–28, 40–40, February 1965. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682329.

Lin:1969:SPB

[LB69] Chi-Yuan Lin and Paul D. Berger. On the selling price and buying price of a lottery. The American Statistician, 23(5):25–26, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682183.

Lerner:1960:FSA

[LCD60] William Lerner, Frank A. Cisar, and Samuel J. Dennis. Federal statistical activities. The American Statistician, 14(4):11–40, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681377.

Lee:1960:SEB

[Lee60] Maurice W. Lee. Statistics and education for business. The American Statistician, 14(1):16–19, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682128.

Lerner:1968:TSP

[Ler68] Monroe Lerner. Training of statistical personnel for state health departments. The American Statistician, 22(4):22–27, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682018.

Lew:1969:LD

[Lew69a] Edward A. Lew. Louis I. Dublin, 1882–1969. The American Statistician, 23(2):33, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681297.

Lew:1969:MS

[Lew69b] Edward A. Lew. Mortimer Spiegelman, 1901–1969. The American Statistician, 23(3):47, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682582.

Li:1964:TAV

[Li64] C. C. Li. Two additional views of linear regression coefficients. The American Statistician, 18(4):27–28, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682154.

Lieberson:1961:NGC

[Lie61] Stanley Lieberson. Non-graphic computation of Kendall’s tau. The American Statistician, 15(4):20–21, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681824.

Litwin:1967:DBE

[Lit67] Samuel Litwin. A decision not based on expected monetary values. The American Statistician, 21(4):25, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682100.

Lindley:1969:LE

[LOG^{+}69] D. V. Lindley, Edward T. O’Donnell, I. J. Good, M. Subramanian, R. Van Wyk, James Pachares, Max L. Eaton, James Stapleton, Bale S. Gurunanjappa, and Richard M. Brugger. Letters to the Editor. The American Statistician, 23(3):48–50, June 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682583.

Levin:1961:ASP

[LS61] Bernard M. Levin and Norman C. Severo. Applications of statistics in post office automation. The American Statistician, 15(4): 14–18, October 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681823.

Lund:1963:SMM

[Lun63] Iver A. Lund. On statistical methods in meteorology. The American Statistician, 17(2):12–13, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2685512.

Lunde:1968:AST

[Lun68] Anders S. Lunde. The Applied Statistics Training Institute (ASTI). The American Statistician, 22(4):17, October 1968. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682016.

Madow:1960:HHT

[Mad60] William G. Madow. Harold Hotelling as a teacher. The American Statistician, 14(3):15–17, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682166.

Mahalanobis:1965:SKT

[Mah65] P. C. Mahalanobis. Statistics as a key technology. The American Statistician, 19(2):43–46, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682378.

Maloney:1962:SAR

[Mal62] Clifford J. Maloney. Statistics in Army research development and testing. The American Statistician, 16(3):13–17, June 1962. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682069.

Mantel:1969:FAV

[Man69a] Nathan Mantel. Functional averages of a variable. The American Statistician, 23(1):21–22, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681685.

Mantel:1969:MLB

[Man69b] Nathan Mantel. More light bulb statistics. The American Statistician, 23(2):21–23, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681293.

Mark:1960:FCC

[Mar60] Mary Louise Mark. Fred C. Croxton, 1871–1960. The American Statistician, 14(3):43, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682173.

Markowitz:1968:MMS

[Mar68a] Etan Markowitz. Minimum mean-square-error estimation of the standard deviation of the normal distribution. The American Statistician, 22(3):26, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor. org/stable/2681801. See comments [Bol68, Cur68b, Bol68] and priority acknowledgement [Mar68b].

Markowitz:1968:PAM

[Mar68b] Etan Markowitz. Priority acknowledgement to “Minimum Mean-Square-Error Estimation of the Standard Deviation of the Normal Distribution”. The American Statistician, 22(4):42, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682027. See

[Mar68a].

McCracken:1969:GPE

[McC69] Paul W. McCracken. The game plan for economic policy. The American Statistician, 23(4):7–10, October 1969. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681724.

McHugh:1961:CII

[McH61] Richard B. McHugh. Confidence interval inference and sample size determination. The American Statistician, 15(2):14–17, April 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682727.

Mathusz:1965:LE

[MDS^{+}65] D. V. Mathusz, Stuart C. Dodd, Joseph N. Skwish, James G. Wendel, Donald E. Gates, and Nura D. Turner. Letters to the Editor. The American Statistician, 19(2):57–58, April 1965. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682385.

Meissner:1960:ASE

[Mei60] Frank Meissner. Adventures in statistics via educational TV. The American Statistician, 14(2):19–20, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682244.

Mendelssohn:1968:CEA

[Men68] Rudolph C. Mendelssohn. The computer and economic analysis at the Bureau of Labor Statistics. The American Statistician, 22 (2):16–19, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681980.

Meyer:1969:LCO

[Mey69] Richard M. Meyer. The k-th largest coordinate of an ordered ntuple. The American Statistician, 23(5):27, December 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682184.

Mantel:1968:WCC

[MG68] Nathan Mantel and Samuel W. Greenhouse. What is the continuity correction? The American Statistician, 22(5):27–30, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681839.

Mihram:1969:CNR

[Mih69] George A. Mihram. A cautionary note regarding invocation of the central limit theorem. The American Statistician, 23(5):38, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682190.

McKenna:1966:LE

[MING66] James F. McKenna, David Ibbetson, Hans Neisser, and N. T. Gridgeman. Letters to the Editor. The American Statistician, 20 (4):41–42, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685965.

Mitra:1969:AAA

[Mit69] S. Mitra. An alternative approach to the analysis of a fourfold table. The American Statistician, 23(5):22–24, December 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682182.

Moore:1966:LE

[MNC^{+}66] Geoffrey H. Moore, Hans Neisser, F. Eugene Clark, Morton J. Netzorg, and Lurgan Sahib. Letters to the Editor. The American Statistician, 20(2):37–38, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682716.

Moore:1967:WR

[Moo67] Geoffrey H. Moore. What is a recession? The American Statistician, 21(4):16–19, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682096.

Moriyama:1966:LJR

[Mor66] Iwao M. Moriyama. Lowell J. Reed, 1886–1966. The American Statistician, 20(3):50, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681503.

Mosteller:1962:CCT

[Mos62] Frederick Mosteller. Continental Classroom’s TV course in probability and statistics. The American Statistician, 16(5):20–25, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681960.

Mosteller:1964:SWS

[Mos64] Frederick Mosteller. Samuel S. Wilks: Statesman of statistics. The American Statistician, 18(2):11–17, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682393.

Mosteller:1967:PRT

[Mos67] Frederick Mosteller. The President reports: Three major ASA actions. The American Statistician, 21(4):2–4, October 1967. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682092.

Murthy:1966:NER

[MP66] V. N. Murthy and C. S. Pillai. A note on the expectation of the reciprocal and square root of a random variable. The American Statistician, 20(5):30, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682640.

Mantel:1968:COP

[MP68] Nathan Mantel and Bernard S. Pasternack. A class of occupancy problems. The American Statistician, 22(2):23–24, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681983.

Minton:1969:FYC

[MS69] Paul D. Minton and Mary Howard Smith. Fifteen years of cooperation: The Southern Regional Committee on Statistics. The American Statistician, 23(4):23–26, October 1969. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681729.

Mullen:1967:NRT

[Mul67] Kenneth Mullen. A note on the ratio of two independent random variables. The American Statistician, 21(3):30–31, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682042.

Mullen:1969:NPC

[Mul69] K. Mullen. A note on particle counting. The American Statistician, 23(4):40–41, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681741.

Murphy:1962:WIK

[Mur62] Thomas E. Murphy. Willford Isbell King. The American Statistician, 16(5):30, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681962.

Magill:1960:GCS

[MWR60] W. A. Magill, Fred. E. Whitworth, and D. L. Ralston. Government of Canada statistical activities. The American Statistician, 14(1):15, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682127.

Natrella:1960:RBC

[Nat60] Mary G. Natrella. The relation between confidence intervals and tests of significance — a teaching aid. The American Statistician, 14(1):20–38, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682129.

Naus:1968:EBP

[Nau68] Joseph I. Naus. An extension of the birthday problem. The American Statistician, 22(1):27–29, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681879. See note [Geh68].

Nelson:1967:EPD

[Nel67] Wayne Nelson. Estimation for a population depleted by sampling. The American Statistician, 21(4):19–21, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682097.

Nemenyi:1969:VEP

[Nem69] Peter Nemenyi. Variances: An elementary proof and a nearly distribution-free test. The American Statistician, 23(5):35–37, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682189.

Nelson:1968:NRT

[NFE68] Norma Nelson, Arnold Feldman, and Lila Elveback. A note on response-to-treatment percentages in cancer. The American Statistician, 22(1):22–23, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681877.

onHistory:1969:APA

[oH69] CBMS Advisory Committee on History. An appeal for preservation of archival materials. The American Statistician, 23(5):17– 18, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682180.

Oster:1962:LE

[Ost62] Gilbert M. Oster. Letter to the Editor. The American Statistician, 16(5):36, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681964.

OToole:1966:PST

[O’T66a] A. L. O’Toole. Probability and statistics teaching in Western European secondary schools. The American Statistician, 20(2): 23–24, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682712.

OToole:1966:SPS

[O’T66b] A. L. O’Toole. Status of probability and statistics in some European universities. The American Statistician, 20(5):22–25, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682638.

Pabst:1960:GB

[Pab60] William R. Pabst, Jr. Grant I. Butterbaugh, 1893–1960. The American Statistician, 14(5):24, December 1960. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682482.

Paroush:1969:IPT

[Par69] Jacob Paroush. Integer programming technique to construct statistical tests. The American Statistician, 23(5):43–44, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682194.

Pierce:1969:IND

[PD69] David A. Pierce and Richard L. Dykstra. Independence and the normal distribution. The American Statistician, 23(4):39, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681740.

Peters:1965:SEF

[Pet65] William S. Peters. Some error formulas when occurrences are known but non-occurrences must be estimated. The American Statistician, 19(5):30–31, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681896.

Pratt:1969:LE

[PGH^{+}69] John W. Pratt, Ernest W. Grove, H. O. Hartley, James E. Grizzle, Salem H. Khamis, and Richard M. Brugger. Letters to the Editor. The American Statistician, 23(2):34–35, April 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681298.

Pachares:1968:LE

[PH68] James Pachares and Peter J. Huber. Letters to the Editor. The American Statistician, 22(4):50, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682030.

Pillai:1968:NEI

[Pil68] C. S. Pillai. A note on the expectation of an irrational function of a random variable. The American Statistician, 22(3):31, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681808. See note

[Kru68].

Pabst:1967:LE

[PKC67] William R. Pabst, Jr., Richard G. Krutchkoff, and Victor Chew. Letters to the Editor. The American Statistician, 21(5):33–35, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682660.

Poch:1960:TSS

[Poc60] F. Azorin Poch. The teaching of statistics in Spain. The American Statistician, 14(4):23–26, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681380.

Porter:1967:EPS

[Por67] Richard C. Porter. Extra-point strategy in football. The American Statistician, 21(5):14–15, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682653.

Pratt:1962:LE

[Pra62] John W. Pratt. Letter to the Editor. The American Statistician, 16(2):24, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681540.

Prince:1967:LTA

[Pri67] Melvin Prince. Life table analysis of prime time programs on a television network. The American Statistician, 21(2):21–23, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682608.

Potthoff:1965:LE

[PWBR65] Richard F. Potthoff, Maurice Whittinghill, Clyde A. Bridger, and David Rothman. Letters to the Editor. The American Statistician, 19(5):41–43, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681900.

Quenouille:1967:LE

[QGS^{+}67] M. H. Quenouille, N. T. Gridgeman, Irving H. Siegel, J. C. Koop, N. L. Johnson, Armand Seri, Richard M. Brugger, David Gomberg, Robert S. Weiss, Thomas A. Goldman, Michael H. Hoyle, Lawrence P. Seidman, Herbert Sichel, Arthur F. Johnson, James A. Hagans, Joseph Berkson, H. W. Norton, and László Palfalvi. Letters to the Editor. The American Statistician, 21(3): 38–44, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682045.

R:1968:SS

[R.68] E. R. The study of statistics. The American Statistician, 22 (1):33–35, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681883.

Rubin:1966:QAb

[RA66] Ernest Rubin and Henry Anderson. Questions and answers. The American Statistician, 20(2):29–33, April 1966. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682714.

Rastogi:1969:NDT

[Ras69] Suresh C. Rastogi. Note on the distribution of a test statistic. The American Statistician, 23(5):40–41, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682192.

Rothman:1965:CAS

[RBB65] Abe Rothman, Robert S. Barker, and Edgar M. Bisgyer. Characteristics of the American Statistical Association membership, 1964. The American Statistician, 19(5):32–36, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681897.

Rangarajan:1969:NCB

[RC69] C. Rangarajan and S. Chatterjee. A note on comparison between correlation coefficients of original and transformed variables. The American Statistician, 23(4):28–29, October 1969. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681732.

Rubin:1963:QAd

[RCF63] Ernest Rubin, Leo G. Connolly, and A. Gerald Fryling. Questions and answers. The American Statistician, 17(4):28–30, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682596.

Rubin:1961:QAb

[RE61] Ernest Rubin and J. Edward Ely. Questions and answers. The American Statistician, 15(2):23–26, April 1961. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682730.

Riley:1960:SDS

[RF60] Donald C. Riley and David L. Futransky. The supply of and demand for statisticians. The American Statistician, 14(4):25– 26, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681381.

Rubin:1969:QAd

[RG69] Ernest Rubin and I. J. Good. Questions and answers. The American Statistician, 23(4):42–45, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681742.

Raff:1966:LE

[RGH66] Morton S. Raff, Ernest Wilson Grove, and Ralph Haertel. Letters to the Editor. The American Statistician, 20(1):34–36, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681677.

Rubin:1961:QAd

[RH61] Ernest Rubin and Ethel D. Hoover. Questions and answers. The American Statistician, 15(4):32–34, October 1961. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681826.

Rubin:1966:QAa

[RH66] Ernest Rubin and John Joseph Hooker. Questions and answers. The American Statistician, 20(1):29–31, February 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681675.

Rice:1964:WFW

[Ric64] Stuart A. Rice. Walter Francis Willcox, March 22, 1861–October 30, 1964. The American Statistician, 18(5):25–26, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682466.

Rice:1967:CGB

[Ric67] Stuart A. Rice. Conception, gestation and birth of IASI. The American Statistician, 21(3):15–19, June 1967. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682038.

Riley:1960:FSA

[RK60] Donald C. Riley and James W. Knowles. Federal statistical activities. The American Statistician, 14(5):10–11, December 1960. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682475.

Rubin:1967:QAe

[RM67] Ernest Rubin and Geoffrey H. Moore. Questions and answers. The American Statistician, 21(5):27–32, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682659.

Rubin:1967:QAa

[RNCE67] Ernest Rubin, A. M. Niessen, Charles A. Church, Jr., and Paul M. Ellner. Questions and answers. The American Statistician, 21 (1):34–36, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681914.

Rubin:1966:QAc

[RNWR66] Ernest Rubin, Joseph I. Naus, Colin White, and R. S. G. Rutherford. Questions and answers. The American Statistician, 20(3): 42–43, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681499.

Roberts:1966:SDO

[Rob66] Harry V. Roberts. Statistical dogma: One response to a challenge. The American Statistician, 20(4):25–27, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685959.

Roberts:1969:ISP

[Rob69] Charles DeWitt Roberts. The infinite stamp problem. The American Statistician, 23(2):9–10, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681288.

Rose:1964:SNR

[Ros64] R. C. Rose. Samarendra Nath Roy, December 11, 1906–July 23, 1964. The American Statistician, 18(5):26–27, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682467.

Ross:1966:LS

[Ros66] Arthur M. Ross. Living with symbols. The American Statistician, 20(3):15–18, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681492.

Rosenberg:1968:MWT

[Ros68] Lloyd Rosenberg. Mean waiting time as a measure of effectiveness. The American Statistician, 22(4):31–34, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682020.

Rothman:1967:PSI

[Rot67] David Rothman. The principle of significant inconsistency: How to profit from dependent events. The American Statistician, 21 (2):23–24, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682609.

Rubin:1963:QAc

[RPG^{+}63] Ernest Rubin, Edward B. Perrin, Bud Goode, Harold E. Young, Earl L. Atwood, Jack Straley Battell, and Al Horowitz. Questions and answers. The American Statistician, 17(3):41–42, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681342.

Rubin:1961:QAc

[RR61] Ernest Rubin and Luba Richter. Questions and answers. The American Statistician, 15(3):24–27, June 1961. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682629.

Rubin:1965:QAc

[RR65] Ernest Rubin and David Rosenblatt. Questions and answers. The American Statistician, 19(3):36–39, June 1965. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681421.

Rubin:1967:QAc

[RR67] Ernest Rubin and David Rosenblatt. Questions and answers. The American Statistician, 21(3):32–37, June 1967. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682044.

Rubin:1960:QAc

[RS60] Ernest Rubin and Emil D. Schell. Questions and answers. The American Statistician, 14(4):27–30, October 1960. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681382.

Rubin:1961:QAe

[RS61] Ernest Rubin and O. M. Sostad. Questions and answers. The American Statistician, 15(5):19–21, December 1961. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681615.

Rubin:1960:P

[Rub60a] Ernest Rubin. Preface. The American Statistician, 14(3):15, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682165.

Rubin:1960:QAa

[Rub60b] Ernest Rubin. Questions and answers. The American Statistician, 14(1):28–31, February 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682132.

Rubin:1960:QAb

[Rub60c] Ernest Rubin. Questions and answers. The American Statistician, 14(2):23–25, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682246.

Rubin:1960:QAd

[Rub60d] Ernest Rubin. Questions and answers. The American Statistician, 14(5):19–21, December 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682480.

Rubin:1961:QAa

[Rub61] Ernest Rubin. Questions and answers. The American Statistician, 15(1):29–30, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682508.

Rubin:1962:QAa

[Rub62a] Ernest Rubin. Questions and answers. The American Statistician, 16(1):22–24, February 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681949.

Rubin:1962:QAb

[Rub62b] Ernest Rubin. Questions and answers. The American Statistician, 16(2):15–17, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681538.

Rubin:1962:QAc

[Rub62c] Ernest Rubin. Questions and answers. The American Statistician, 16(3):37–39, June 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682073.

Rubin:1962:QAd

[Rub62d] Ernest Rubin. Questions and answers. The American Statistician, 16(4):27–30, October 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681435.

Rubin:1962:QAe

[Rub62e] Ernest Rubin. Questions and answers. The American Statistician, 16(5):27–29, December 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681961.

Rubin:1963:QAa

[Rub63a] Ernest Rubin. Questions and answers. The American Statistician, 17(1):25–26, February 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682493.

Rubin:1963:QAb

[Rub63b] Ernest Rubin. Questions and answers. The American Statistician, 17(2):19–24, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685515.

Rubin:1963:QAe

[Rub63c] Ernest Rubin. Questions and answers. The American Statistician, 17(5):22–24, December 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685673.

Rubin:1964:QAa

[Rub64a] Ernest Rubin. Questions and answers. The American Statistician, 18(2):27–29, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682399.

Rubin:1964:QAb

[Rub64b] Ernest Rubin. Questions and answers. The American Statistician, 18(3):29–32, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681483.

Rubin:1964:QAc

[Rub64c] Ernest Rubin. Questions and answers. The American Statistician, 18(4):38–40, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682156.

Rubin:1964:QAd

[Rub64d] Ernest Rubin. Questions and answers. The American Statistician, 18(5):22–24, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682465.

Rubin:1965:QAa

[Rub65a] Ernest Rubin. Questions and answers. The American Statistician, 19(1):33–35, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682331.

Rubin:1965:QAb

[Rub65b] Ernest Rubin. Questions and answers. The American Statistician, 19(2):49–51, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682380.

Rubin:1965:QAd

[Rub65c] Ernest Rubin. Questions and answers. The American Statistician, 19(4):33–36, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681574.

Rubin:1965:QAe

[Rub65d] Ernest Rubin. Questions and answers. The American Statistician, 19(5):37–40, December 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681898.

Rubin:1966:QAd

[Rub66a] Ernest Rubin. Questions and answers. The American Statistician, 20(4):32–34, October 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685963.

Rubin:1966:QAe

[Rub66b] Ernest Rubin. Questions and answers. The American Statistician, 20(5):32–34, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682642.

Rubin:1967:QAb

[Rub67a] Ernest Rubin. Questions and answers. The American Statistician, 21(2):35–38, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682614.

Rubin:1967:QAd

[Rub67b] Ernest Rubin. Questions and answers. The American Statistician, 21(4):45–48, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682107.

Rubin:1968:QAa

[Rub68a] Ernest Rubin. Questions and answers. The American Statistician, 22(1):31–33, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681882.

Rubin:1968:QAb

[Rub68b] Ernest Rubin. Questions and answers. The American Statistician, 22(2):31–33, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681989.

Rubin:1968:QAc

[Rub68c] Ernest Rubin. Questions and answers. The American Statistician, 22(3):42–43, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681810.

Rubin:1968:QAd

[Rub68d] Ernest Rubin. Questions and answers. The American Statistician, 22(4):43–45, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682028.

Rubin:1968:QAe

[Rub68e] Ernest Rubin. Questions and answers. The American Statistician, 22(5):33–35, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681842.

Rubin:1968:SPK

[Rub68f] Ernest Rubin. Statistics and the perimeters of knowledge. The American Statistician, 22(5):35–36, December 1968. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681843.

Rubin:1969:QAa

[Rub69a] Ernest Rubin. Questions and answers. The American Statistician, 23(1):32–34, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681691.

Rubin:1969:QAb

[Rub69b] Ernest Rubin. Questions and answers. The American Statistician, 23(2):25–30, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681295.

Rubin:1969:QAc

[Rub69c] Ernest Rubin. Questions and answers. The American Statistician, 23(3):43–46, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682580.

Rubin:1969:QAe

[Rub69d] Ernest Rubin. Questions and answers. The American Statistician, 23(5):46–48, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682196.

Sakuntala:1969:SSC

[Sak69] Velpuri Sakuntala. Simulation studies for comparison of different estimators. The American Statistician, 23(5):31–32, December 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682187.

Samuelson:1960:HHM

[Sam60] Paul A. Samuelson. Harold Hotelling as mathematical economist. The American Statistician, 14(3):21–25, June 1960. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682168.

Sandelius:1967:NVD

[San67] Martin Sandelius. A note on the variance of a discrete uniform distribution. The American Statistician, 21(5):21, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682656.

Sasso:1961:TSC

[Sas61] Roberto Sasso. The teaching of statistics in Costa Rica. The American Statistician, 15(2):20–22, April 1961. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682729.

Shaw:1965:RCA

[SBB^{+}65] William H. Shaw, Murray R. Benedict, David H. Blackwell, Lester R. Frankel, Philip H. Hauser, Arnold J. King, Kenneth E. Miller, Geoffrey H. Moore, Marc Nerlove, Chester Rapkin, Lazare Teper, and Ralph J. Watkins. Report of the Census Advisory Committee of the American Statistical Association. The American Statistician, 19(1):22, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682327.

Shiskin:1969:RPC

[SBB^{+}69] Julius Shiskin, Z. W. Birnbaum, Ralph A. Bradley, Herbert A. David, Murray F. Foss, J. Stuart Hunter, John W. Lehman, Fred C. Leone, Margaret E. Martin, John W. Pratt, Ernest Rubin, and David L. Wallace. 1968 report of the publications committee. The American Statistician, 23(2):36–37, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681299.

Stringham:1960:FSA

[SBFC60] Luther Stringham, Gertrude Bancroft, M. Douglas Fahey, and Edward T. Crowder. Federal statistical activities. The American Statistician, 14(3):11–13, June 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682164.

Simon:1966:LE

[SBSK66] Julian L. Simon, Donald C. Butler, John H. Smith, and Richard G. Krutchkoff. Letters to the Editor. The American Statistician, 20 (5):35–37, December 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682643.

Schmitt:1961:SNS

[Sch61] Robert C. Schmitt. Statehood and national statistics. The American Statistician, 15(1):27–30, February 1961. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682507.

Schlaifer:1963:LE

[Sch63a] Robert Schlaifer. Letter to the Editor. The American Statistician, 17(2):36, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685519.

Schleifer:1963:NBS

[Sch63b] Arthur Schleifer, Jr. The new business statistics: Its implementation in the basic required course. The American Statistician, 17 (3):18–23, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681337.

Schor:1967:SRP

[Sch67] Stanley Schor. Statistical reviewing program for medical manuscripts. The American Statistician, 21(1):28–31, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681911.

Schmitt:1968:MSI

[Sch68] Robert C. Schmitt. Migration statistics in an island state: The Hawaii experience. The American Statistician, 22(2):20–23, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681982.

Searls:1967:NUA

[Sea67] Donald T. Searls. A note on the use of an approximately known coefficient of variation. The American Statistician, 21(3):20–21, June 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682039.

Searle:1969:CBM

[Sea69] S. R. Searle. Correlation between means of parts and wholes. The American Statistician, 23(2):23–24, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681294.

Semon:1964:MAD

[Sem64] Thomas T. Semon. Minimizing absolute deviations. The American Statistician, 18(2):25, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682397.

Serfling:1967:NCR

[Ser67] R. J. Serfling. A note on the covariance of a random variable and its reciprocal. The American Statistician, 21(4):33, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682104.

Stanley:1969:APS

[SG69] Julian C. Stanley and Gene V. Glass. An algebraic proof that the sum of the squared errors in estimating Y from X via b_{1} and b_{0} is minimal. The American Statistician, 23(1):25–26, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681687.

Smith:1968:SMO

[SH68] William B. Smith and R. R. Hocking. A simple method for obtaining the information matrix for a multivariate normal distribution. The American Statistician, 22(1):18–20, February 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681874.

Shaw:1967:MBW

[Sha67] Lawrence H. Shaw. Mr. Bean and the weather. The American Statistician, 21(5):12–13, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682651. See rejoinder [Bea67b].

Shah:1969:MPF

[Sha69] S. M. Shah. On the minimum property of the first absolute moment. The American Statistician, 23(3):27, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682577.

Shiskin:1960:HA

[Shi60] Julius Shiskin. How accurate? The American Statistician, 14 (4):15–17, October 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681378.

Shultz:1969:LSN

[Shu69] George P. Shultz. Labor statistics for national decision making. The American Statistician, 23(4):11–14, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681725.

Sibley:1969:SAR

[Sib69] Elbridge Sibley. Stuart Arthur Rice, 1889–1969. The American Statistician, 23(4):47–48, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681744.

Siegel:1968:SAL

[Sie68] Irving H. Siegel. Supermatrix approach to least-squares adjustment. The American Statistician, 22(5):22–23, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681835.

Sealy:1968:FR

[SK68] Elbridge Sealy and Clyde V. Kiser. Frank A. Ross, 1888–1968. The American Statistician, 22(3):14, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681793.

Sanders:1964:LE

[SLG^{+}64] Paul G. Sanders, James A. Lechner, L. V. Goodman, William L. Hafley, J. Hunter McDowell, H. W. Norton, and M. T. L. Bizley. Letters to the Editor. The American Statistician, 18(2):30–31, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682400.

Smith:1963:FSA

[Smi63] Mapheus Smith. Federal statistical activities. The American Statistician, 17(3):13–42, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681335.

Smillie:1964:RRP

[Smi64a] K. W. Smillie. Remarks on a recent paper on round-off errors in regression analysis. The American Statistician, 18(4):26–40, October 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682153.

Smith:1964:LFF

[Smi64b] George Cline Smith. The law of forecast feedback. The American Statistician, 18(5):11–14, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682461.

Smith:1964:STW

[Smi64c] H. Smith. Statistical training within an industrial firm. The American Statistician, 18(2):22–24, April 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682396.

Smith:1968:CS

[Smi68a] John H. Smith. Careers in statistics. The American Statistician, 22(2):34, April 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681990.

Smith:1968:SPP

[Smi68b] Armand V. Smith, Jr. Some probability problems in the game of “Craps”. The American Statistician, 22(3):29–30, June 1968. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681806.

Smith:1969:AFB

[Smi69] H. Fairfield Smith. Approximate formulae for bias and variance of ratios. The American Statistician, 23(4):29–31, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681733.

Somberg:1967:SMS

[Som67] Seymour I. Somberg. A simplified method of solving simple and multiple regression equations. The American Statistician, 21(1): 31–32, February 1967. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681912.

Sprott:1969:NCO

[Spr69] D. A. Sprott. A note on a class of occupancy problems. The American Statistician, 23(2):12–13, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681290.

Selvin:1966:DDP

[SS66] Hanan C. Selvin and Alan Stuart. Data-dredging procedures in survey analysis. The American Statistician, 20(3):20–23, June 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681493.

Sclove:1967:FRE

[SSV67] Stanley L. Sclove, Gordon Simons, and J. Van Ryzin. Further remarks on the expectation of the reciprocal of a positive random variable. The American Statistician, 21(4):33–34, October 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682105.

Stekler:1960:SL

[Ste60a] Herman O. Stekler. On smoothing and lags. The American Statistician, 14(5):13, December 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682477.

Stevenson:1960:RHC

[Ste60b] John A. Stevenson. Robert Hamilton Coats, 1874–1960. The American Statistician, 14(2):28–29, April 1960. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682248.

Stuart:1963:CSR

[Stu63] Alan Stuart. Calculation of Spearman’s rho for ordered two-way classifications. The American Statistician, 17(4):23–24, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682594.

Stuart:1969:RMS

[Stu69] Alan Stuart. Reduced mean-square-error estimation of σ^{p} in normal samples. The American Statistician, 23(4):27–28, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681731.

Subramanian:1963:PSG

[Sub63] S. Subramanian. The present state of government statistics in India. The American Statistician, 17(3):31–40, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681340.

Taeuber:1963:MJH

[Tae63] Conrad Taeuber. Margaret Jarman Hagood, 1908–1963. The American Statistician, 17(4):37, October 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682599.

Taeuber:1964:PKW

[Tae64] Conrad Taeuber. P. K. Whelpton. The American Statistician, 18 (3):20, June 1964. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681479.

Tanner:1960:FGS

[Tan60] William F. Tanner. Filtering in geological sampling. The American Statistician, 14(5):12–21, December 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682476.

Tarter:1967:PIS

[Tar67] Michael Tarter. Programmed instruction in statistics from the professor’s point of view. The American Statistician, 21(2):28–31, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682612.

Taylor:1965:LE

[Tay65] Kenneth D. Taylor. Letters to the Editor. The American Statistician, 19(1):16, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682325.

Taylor:1966:RIB

[Tay66] William F. Taylor. A report on an inventory of biometry. The American Statistician, 20(5):16–20, December 1966. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682637.

Thompson:1961:RSS

[TE61] W. A. Thompson, Jr. and Jack Endriss. The required sample size when estimating variances. The American Statistician, 15(3): 22–23, June 1961. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682628.

Terry:1968:CLM

[Ter68] Edwin F. Terry. Class limits and midpoints. The American Statistician, 22(3):28, June 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681804.

Thompson:1967:EP

[Tho67] W. A. Thompson, Jr. Economic probability. The American Statistician, 21(2):26–27, April 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682611.

Taeuber:1967:SSR

[TMW67] Conrad Taeuber, Frederick Mosteller, and Paul Webbink. Social Science Research Council Committee on Statistical Training. The American Statistician, 21(5):10–11, December 1967. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682650.

Tukey:1965:TTSb

[Tuk65] John W. Tukey. The technical tools of statistics. The American Statistician, 19(2):23–28, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682374.

Turner:1962:EDS

[Tur62] Nura D. Turner. The expected dependency of students. The American Statistician, 16(2):14, April 1962. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681537.

Ury:1963:LE

[Ury63] Hans Ury. Letter to the Editor. The American Statistician, 17 (3):54, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681350.

VanderNoot:1968:CSS

[Van68] T. J. Vander Noot. Computers, social science statisticians, and the A.S.A. The American Statistician, 22(3):20–22, June 1968. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681798.

Valentine:1969:LE

[VKT^{+}69] Thomas J. Valentine, Salem H. Khamis, C. J. Thomas, Francis Emeric Binet, R. C. Juola, John H. Smith, I. J. Good, W. Edwards Deming, James A. Merrill, Samuel H. Plotkin, A. L. Wilson, and Jack Tomsky. Letters to the Editor. The American Statistician, 23(4):51–55, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2681747.

Wallis:1965:NYA

[Wal65a] W. Allen Wallis. The next 125 years of the American Statistical Association. The American Statistician, 19(2):39–43, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682377.

Wallis:1965:PCa

[Wal65b] W. Allen Wallis. The President’s column. The American Statistician, 19(1):2, February 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682319.

Wallis:1965:PCb

[Wal65c] W. Allen Wallis. The President’s column. The American Statistician, 19(2):2, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682367.

Ward:1969:SRM

[War69] Joe H. Ward, Jr. Synthesizing regression models — an aid to learning effective problem analysis. The American Statistician, 23(2):14–20, April 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681292.

Wasserman:1963:FSG

[Was63] William Wasserman. Foundations of statistics and goals of the basic required course. The American Statistician, 17(3):15–18, June 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681336.

Watts:1965:TGP

[Wat65] Harold W. Watts. The Test-O-Gram: a pedagogical and presentational device. The American Statistician, 19(4):22–28, October 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681570.

Watkins:1960:DD

[WCC^{+}60] Ralph J. Watkins, William J. Carson, Gertrude M. Cox, Martin R. Gainsbrugh, Walter E. Hoadley, Jr., Howard L. Jones, William R. Leonard, Rensis Likert, Isador Lubin, Frederick F. Stephan, Willard L. Thorp, and Samuel S. Wilks. E. Dana Durand. The American Statistician, 14(2):29, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682249.

Wilson:1969:NNL

[WD69] A. L. Wilson and A. W. Douglas. A note on non-linear curve fitting. The American Statistician, 23(4):37–38, October 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681739.

Woolsey:1960:FSA

[WDC^{+}60] Theodore D. Woolsey, Samuel J. Dennis, Hugh Carter, Thomas J. Mills, Raymond Nassimbene, and Helen F. Demond. Federal statistical activities. The American Statistician, 14(2):8–25, April 1960. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682241.

Webb:1965:DMS

[Web65] James E. Webb. Decision-making and statistical standards. The American Statistician, 19(5):16–20, December 1965. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681893.

Wernimont:1966:RAC

[Wer66] Grant Wernimont. Report to ASTM Committee E-11 of round table discussion on statistics in the testing of materials. The American Statistician, 20(2):42–43, April 1966. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682719.

Wolfbein:1961:FSA

[WGW^{+}61] Seymour L. Wolfbein, Edwin D. Goldfield, Robert M. Walsh, Max Schiferl, and Saul Hern. Federal statistical activities. The American Statistician, 15(3):11–13, June 1961. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2682623.

White:1964:UCS

[Whi64] Colin White. Unkind cuts at statisticians. The American Statistician, 18(5):15–17, December 1964. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682462.

Wilson:1965:STN

[Wil65] Edwin B. Wilson. Statistics, then and now. The American Statistician, 19(2):37–38, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682376.

Wilkinson:1968:WSP

[Wil68] Bryan Wilkinson. World series pools. The American Statistician, 22(4):35, October 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682021.

Wisniewski:1966:ASS

[Wis66] T. K. M. Wisniewski. Another statistical solution of a combinatorial problem. The American Statistician, 20(3):25, June 1966. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681495.

Willcox:1961:WFW

[WL61] Walter F. Willcox and William R. Leonard. Walter F. Willcox: Statist. The American Statistician, 15(1):16–19, February 1961. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682503.

Wolfbein:1965:NCW

[Wol65] Seymour L. Wolfbein. On the nature and conditions of working life. The American Statistician, 19(2):19–23, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682373.

Worcester:1965:EBW

[Wor65] Jane Worcester. Edwin Bidwell Wilson. The American Statistician, 19(2):52–53, April 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682381.

Working:1968:JW

[Wor68] Mrs. E. J. Working. E. J. Working, 1900–1968. The American Statistician, 22(5):37, December 1968. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681844.

Walls:1969:NVP

[WW69] Robert C. Walls and David L. Weeks. A note on the variance of a predicted response in regression. The American Statistician, 23 (3):24–26, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682576.

Young:1963:GSE

[You63] Dudley E. Young. Governmental statistics — an example of State– Federal cooperation. The American Statistician, 17(2):16–18, April 1963. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685514.

Zarnowitz:1969:NAN

[Zar69] Victor Zarnowitz. The new ASA–NBER survey of forecasts by economic statisticians. The American Statistician, 23(1):12–16, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681682.

Zelen:1969:EB

[Zel69] Marvin Zelen. The education of biometricians. The American Statistician, 23(4):14–15, October 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2681726.